### v1

# Duplex-GS: Proxy-Guided Weighted Blending for Real-Time Order-Independent Gaussian Splatting

Abstract
Recent advances in 3D Gaussian Splatting (3DGS) have demonstrated remarkable rendering fidelity and efficiency. However, these methods still rely on computationally expensive sequential $\alpha$-blending operations, resulting in significant overhead—particularly on resource-constrained platforms. In this paper, we propose Duplex-GS, a dual-hierarchy framework that integrates proxy Gaussian representations with order-independent rendering techniques to achieve photorealistic results while sustaining real-time performance. To mitigate the overhead caused by view-adaptive radix sort, we introduce cell proxies for local Gaussians management and propose cell search rasterization for further acceleration. By seamlessly combining our framework with Order-Independent Transparency (OIT), we develop a physically inspired weighted sum rendering technique that simultaneously eliminates “popping” and “transparency” artifacts, yielding substantial improvements in both accuracy and efficiency. Extensive experiments on a variety of real-world datasets demonstrate the robustness of our method across diverse scenarios, including multi-scale training views and large-scale environments. Our results validate the advantages of the OIT rendering paradigm in Gaussian Splatting, achieving high-quality rendering with an impressive $1.5\times$$4\times$ speedup over existing OIT based Gaussian Splatting approaches and $52.2\%\,$$86.9\%$ reduction of the radix sort overhead without quality degradation.
Index Terms:
Novel View Synthesis, Gaussian Splatting, Order-Independent Transparency, Real-Time Rendering
I Introduction
Novel view synthesis (NVS) and real-time rendering have attracted significant attention in both academia and industry, with broad applications in video games, virtual reality (VR) [1], and 3D scene reconstruction [2, 3, 4, 5]. Inspired by the groundbreaking Neural Radiance Field (NeRF) [6], learning based methods have substantially advanced the realism of digital content, narrowing the gap between the virtual and real worlds. However, these methods often entail considerable computational overhead. The emergence of 3D Gaussian Splatting (3DGS) [7, 8] marks another milestone, enabling photorealistic and real-time rendering of radiance fields via rasterization—a paradigm akin to the graphics pipeline in modern graphics processing units (GPUs).
Despite the impressive reconstruction and rendering capabilities of 3DGS, a primary limitation is the need for an ever-increasing number of Gaussian primitives to fit all training views. This results in substantial and irregular memory access during rendering, as costly view-dependent sorting operations over long sequences impose considerable computational and memory overhead [9, 10]. Moreover, the sorting process can introduce visible “popping” artifacts [11], as illustrated in Fig. 1. Collectively, these limitations hinder the scalability of 3DGS for large-scale applications and restrict its deployment on resource-constrained platforms such as mobile phones and VR devices.
More recently, sort-free Gaussian Splatting [12] has incorporated Order-Independent Transparency (OIT) [13, 14] into the Gaussian Splatting framework, eliminating the need for sorting in the rasterization pipeline by employing Weighted Sum Rendering (WSR) [15]. This effectively eliminates “popping” artifacts and has been successfully demonstrated on mobile devices [12]. However, as illustrated in Fig. 1, the absence of early termination when encountering opaque objects leads to “transparency” artifacts and increased rendering time, particularly in large-scale scenes, due to the accumulation of redundant Gaussian primitives in final blending.
Neural Gaussian framework [16, 17] is another line of work to enhance storage efficiency for 3DGS, which transitions from explicit 3D Gaussian primitives to a hybrid representation where clusters of Gaussians are encoded via “anchors” and associated with a shared MLP. However, this paradigm adheres to the conventional rendering pipeline and still operates on sets of 3D Gaussians decoded on-the-fly by the MLP. As a result, sorting the lengthy sequence of Gaussians remains mandatory for such Tile based Rendering (TBR), presenting substantial challenges for devices with limited resources [10]. Furthermore, since the spatial positions of neural Gaussians are not explicitly constrained by the anchor points, the generated Gaussians may deviate significantly from their associated anchors—as illustrated in Fig. 2—thereby complicating anchor management.
Figure 1: Motivation of Duplex-GS. The rendering process for 3DGS rely on sorting for long tile-Gaussian sequence, introducing popping artifacts. The sort-free LC-WSR uses weighted sum rendering method without the need of consuming sorting stage, introducing transparency artifacts due to the lack of early termination when encounter opaque object.
Figure 2: Overview of the proposed Duplex-GS. The dual-hierarchy structure introduce cell as the proxy of local similar Gaussians. (a) Compared with proxy anchor, the proposed proxy cell is an ellipsoid with specific geometry. (b) Combing OIT rendering with 3DGS models, our method successfully remove the “popping” and “transparency” artifacts simultaneously. (c) By introducing cell search rasterization (Sec. IV) and physical weighted sum rendering (Sec. V), we achieve significant improvement over OIT Gaussian Splatting for both accuracy and efficiency.
In this paper, we present Duplex-GS, which models 3D scenes with WSR rendering paradigm and dual-layered structure following the neural Gaussian framework. Firstly, scenes are encoded as explicit proxies that are easily stored and manipulated in the first layer. As the concept anchor degenerates the 3D Gaussian to point cloud, we introduce “cell” as the replacement, which is an ellipsoid with concrete physical geometry initialized at the Structure-from-Motion (SfM) [18] points. By constraining the emitted neural Gaussians located inside the cell region, the rasterization can be performed in cell level. Secondly, primitives are decoded as other forms, which is suitable for high-fidelity rendering in the second layer. Encouraged by [12], we introduce the physical inspired WSR with early termination where the specific order of Gaussian depth is not required, which replaces the non-commutative blending operations with parallelizable weighted sum operation. As the number of primitives in the first layer is extensively fewer than that of Gaussians, the proposed cell search rasterization significantly reduce the memory and computational overhead introduced by radix sort. On the one hand, rasterization and lightweight sorting in the first layer removes the “transparency” artifact. On the other hand, the WSR kernel removes the “popping” artifact introduce by vanilla $\alpha$-blending.
In summary, the contributions of this work are as follow:
  • We introduce a dual-layered Gaussian hierarchy, in which the first layer consists of proxies called cells and the second layer comprises 3D Gaussians. Our approach significantly alleviates the need for global sorting while preserving the parallelizability of WSR, thereby achieving artifact-free, real-time 3D reconstruction.
  • Cell rasterization in layer I: We introduce cells—ellipsoids with adaptive geometric properties—to organize emitted Gaussians. Based on this structure, we propose a novel cell rasterization strategy that dramatically reduces both memory and computational overhead associated with radix sort.
  • Physical inspired WSR blending in layer II: We retrieve physical constraints of Gaussian parameters for WSR method to blend Gaussians without the need for global sorting, inherently avoiding the “popping” artifact. Building on this, we introduce cell-level early termination to further accelerate rendering and effectively mitigate the “transparency” artifact encountered in the original WSR approach.
  • We implement WSR based Gaussian Splatting on CUDA-enabled platforms. Experimental results demonstrate that our method delivers comparable rendering speed and visual fidelity to existing approaches on commercial high-end GPUs.
II Related Work
II-A Neural Representation for 3D Scene
Traditional scene reconstruction methods, such as SfM [18, 19] and Multi-View Stereo (MVS) [20, 21], often struggle to recover texture-less regions or handle challenging lighting variations. The advent of Neural Radiance Fields (NeRF) [6] has demonstrated the impressive capability of learning based models to achieve photorealistic reconstructions from sparse 2D images. Building on this, numerous efficient NeRF variants have been developed [22, 23, 24], making real-time rendering increasingly feasible on edge devices [25]. A significant breakthrough in the field is 3D Gaussian Splatting (3DGS) [7], which leverages explicit anisotropic 3D Gaussians for scene reconstruction and substantially accelerates rendering. Nevertheless, the substantial computational and memory demands of 3DGS continue to limit its applicability on resource-constrained platforms. Notably, a trend is witnessed that as scene representations become more explicit, rendering speed improves at the cost of increased memory requirements. And hybrid variants have emerged as a promising compromise between rendering efficiency and memory usage [22, 26, 16]. For instance, Instant-NGP [22] employs an explicit learnable hash table that maps 3D space to a multi-resolution grid, thereby alleviating the computational burden on the MLP. Scaffold-GS [16] introduces a hierarchical structure that encodes neural Gaussians using anchor points, achieving comparable rendering quality and speed with a significantly reduced model size. These advances enable fast, high-resolution rendering without compromising visual fidelity. However, these approaches still rely on the conventional blending pipeline requiring the Gaussians to be rendered in sorted depth order, typically achieved with tile based sorting algorithms. This sorting step not only complicates practical implementation but also introduces visual artifacts such as sudden changes in appearance, commonly referred to as “popping” artifacts [11].
II-B Efficient Neural Rendering
Unlike NeRF whose constraint is the high computational demand, the bottlenecks of 3DGS practical deployment are the extensive number of Gaussian primitives and the frequent memory access requirements [10, 9]. Various works aim to improve the efficiency of 3DGS by reducing the model size or the number of Gaussians involved in the sorting based blending operations [17, 27, 28, 29]. Encouraged by the OIT [13, 14], a widely deployed rendering paradigm in modern GPU to blend non-opaque media, sort-free Gaussian Splatting [12] introduce a novel blending paradigm WSR which eliminates the need of sorting operation in 3DGS, which removes the “popping” artifact as well as achieves faster rendering speed on mobile devices with small scale scenes. However, the lack of early termination results in the “transparency” artifacts as well as introduces extra computation when encounter opaque objects.
II-C Order Independent Transparency
Order-Independent Transparency (OIT) [13, 14] is a foundational graphics technique integrated into modern GPUs to enable the correct compositing of transparent fragments without requiring explicit depth sorting prior to rasterization. Early approaches, such as Depth Peeling [30], A-buffer [31], and Multi-Layer Alpha Blending (MLAB) [32], successfully achieve OIT but often incur significant time or memory overhead. Weighted Blended OIT [15] offers a more efficient alternative by approximating the integration of transparency in a single render pass, storing weighted sums of color and opacity with carefully designed blending weights. And this method has been applied to 3D Gaussian Splatting (3DGS) rendering [12]. However, the absence of early termination in [12] when encountering opaque regions can degrade both visual quality and rendering performance. In contrast, the proposed Duplex-GS framework integrates the core idea of early termination from depth peeling—halting compositing upon encountering fully opaque surfaces—with Weighted Blended OIT. As a result, Duplex-GS attains rendering quality and speed on par with state-of-the-art (SOTA) 3DGS models on high-end GPUs, while exhibiting strong potential for deployment on resource-constrained edge devices by alleviating the huge overhead of the view-adaptive sorting stage.
Figure 3: Illustration of the proposed Duplex-GS pipeline. A dual-layered structure is designed to organize cells and Gaussians. In Layer I, proxy cells—represented as ellipsoids with specific geometric properties—are designed to aggregate Gaussians, enabling rasterization at the cell level and thereby alleviating the overhead associated with view-adaptive global sorting. Leveraging shared MLPs, the cell primitives are subsequently decoded into Gaussian primitives. In Layer II, the decoded Gaussians are blended using weighted sum rendering. Additionally, early termination technique is implemented—guided by the sorting results of the sparse cells from Layer I—which facilitates real-time, photorealistic rendering.
III Preliminaries
III-A Gaussian Splatting
3D Gaussian Splatting [7] models scenes using anisotropic 3D Gaussian primitives and renders images through a rasterization based paradigm. Each 3D Gaussian $G(\bm{x})$ is represented by a center coordinate $\bm{\mu}\in\mathbb{R}^{3}$ and a covariance matrix $\bm{\Sigma}\in\mathbb{R}^{3\times 3}$:
$G(\bm{x})=e^{-\frac{1}{2}(\bm{x}-\bm{\mu})^{T}\bm{\Sigma}^{-1}(\bm{x}-\bm{\mu})},$ (1)
where $\bm{x}\in\mathbb{R}^{3}$ is the queried position of the Gaussian kernel, and the covariance $\bm{\Sigma}$ is parameterized by scales $\bm{s}\in\mathbb{R}^{3}$ and quaternions $\bm{q}\in\mathbb{R}^{4}$. In line with volumetric rendering conventions, each Gaussian is further associated with an opacity $\sigma\in\mathbb{R}$ and a view-dependent color $\bm{c}\in\mathbb{R}^{3}$, which is commonly encoded using Spherical Harmonics (SH). Rendering proceeds via a tile based Gaussian rasterization technique, which employs efficient radix sort to enable $\alpha$-blending
$\bm{C}(\bm{x}^{\prime})=\sum_{i\in N}T_{i}^{\text{3DGS}}\bm{c}_{i}\alpha_{i},\quad\alpha_{i}=\sigma_{i}G^{\prime}_{i}(\bm{x}^{\prime}),$ (2)
where $\bm{x}^{\prime}$ denotes the 2D coordinates of the queried pixel, $G^{\prime}_{i}(\bm{x}^{\prime})$ is the projected 2D gaussian kernel, and $T_{i}^{\text{3DGS}}$ denotes the transmittance
$T_{i}^{\text{3DGS}}=\prod_{j=1}^{i-1}(1-\alpha_{i}).$ (3)
III-B Scaffold-GS
Scaffold-GS introduces anchors as local encoding proxies, which emit neural Gaussians decoded from latent features to effectively represent local scene structure. Specifically, each anchor emits a predefined set of $K$ neural gaussians, whose centers are given by:
$\bm{\mu}_{n,k}=\bm{x}_{n}+\mathcal{O}_{n,k}^{\text{pos}}\cdot\bm{S}_{n},$ (4)
where $n\in\{1,2,...,N\}$ indexes the anchors, $k\in\{1,2,...K\}$ indexes the neural Gaussians per anchor, $\bm{x}_{n}$ denotes the center of anchor-$n$, and $\bm{\mu}_{n,k}$ represents the center of the emitted neural Gaussian-$k$. $\mathcal{O}_{n,k}^{\text{pos}}$ is the learned offset scaled by $\bm{S}_{n}$. Additionally, the opacities of the $K$ neural Gaussians are decoded from the latent feature $\bm{f}_{n}$, viewing distance $\delta_{n}$ and direction $\bm{r}_{n}$ via a scene-wise MLP
$\{\sigma_{n,1},...,\sigma_{n,K}\}=F_{\sigma}(\bm{f}_{n},\delta_{n},\bm{r}_{n}),$ (5)
where $F_{\sigma}$ is the shared opacity decoder. Similarly, the other attributes of each neural Gaussians, including color $\bm{c}_{n,k}$, scales $\bm{s}_{n,k}$ and quaternions $\bm{q}_{n,k}$, are predicted by the separate MLPs $F_{c}$, $F_{s}$, $F_{q}$, respectively.
III-C Weighted Sum Rendering
Inspired by Order Independent Transparency (OIT)—a widely adopted technique for rendering transparent media in traditional graphics pipelines—sort-free Gaussian Splatting integrates this concept into 3DGS, producing the final image as
$\bm{C}=\frac{\bm{c}_{B}w_{B}+\sum_{i=1}^{N}\bm{c}_{i}\alpha_{i}w(d_{i})}{w_{b}+\sum_{i=1}^{N}\alpha_{i}w(d_{i})},$ (6)
where $\bm{c}_{B}$ and $w_{B}$ denote the background color and the learnable background weights respectively, and $d_{i}$ indicates the depths of the $i$-th 3D Gaussians.
By employing this WSR formulation, the non-commutative, computationally intensive sorting stage—traditionally required for compositing—is eliminated. This significantly enhances the potential for parallelization and accelerates rendering. Specifically, the SOTA Linear Correction weighted-sum rendering (LC-WSR) function $w(d_{i})$ is defined as
$w(d_{i})=T_{i}^{\text{LC-WSR}}\cdot v_{i},$ (7)
$T_{i}^{\text{LC-WSR}}=\max\left(0,1-\frac{d_{i}}{\tau}\right),$ (8)
where $\tau$ and $v_{i}$ are the learnable parameters. This function formulates the decay trend w.r.t. the depths of the rendering Gaussians.
IV Layer I: Cell Rasterization for 3DGS
While neural Gaussian frameworks [16, 17] have achieved remarkable reconstruction accuracy, the use of anchors as encoding proxies lacks a concrete geometric interpretation. So we redesign the neural Gaussian proxies to alleviate the computational overhead associated with sorting and rasterization. Specifically, instead of encoding 3D Gaussians with anchors, we introduce the concept of “cell”—an ellipsoid with explicit geometric properties and a defined visible region—to encode neural Gaussians together with their latent features. Based on this representation, we further propose a cell rasterization strategy, enabling efficient and lightweight rasterization.
IV-A Cell Proxy
Unlike anchors, which are points lacking geometric structure, each cell is represented as an ellipsoid that contains several neural 3D Gaussians whose physical properties are dynamically decoded from a trainable feature.
Initialization begins with SfM points $\mathbf{P}\in\mathbb{R}^{N\times 3}$, from which we generate the initial cells, each centered at a point in accordance with the method described in [17]. Every cell is parameterized by physical structure scales $\bm{S}_{n}$ and and quaternions $\bm{Q}_{n}$, defining its visible region. In the dual-layered hierarchy, each cell is further associated with a latent feature vector $\bm{f}_{n}$ that encodes local scene information, and emits $K$ neural Gaussians as described in Eq. 4. The attributes of these $K$ neural Gaussians—including color $\bm{c}_{n,k}$, scales $\bm{s}_{n,k}$ and quaternions $\bm{q}_{n,k}$—are decoded from the latent feature $\bm{f}_{n}$, along with the viewing distance $\delta_{n}$ and viewing direction $\bm{d}_{n}$, via shared scene-wise MLPs.
To ensure that each Gaussian center lies within the visible region of its corresponding cell, we constrain the offset parameter $\mathcal{O}_{n,k}^{\text{pos}}$ to the range [1,1][-1,1]. Additionally, to align the collective regions of the emitted neural Gaussians with the visible extent of the cell, we modify the encoding of the scale $\bm{s}_{n,k}$ and quaternions $\bm{q}_{n,k}$ attributes, thereby directly linking the geometric properties of the cell and its Gaussians
$\{\mathcal{O}^{\text{scale}}_{n,1},...,\mathcal{O}^{\text{scale}}_{n,K}\}=F_{s}(f_{n},\delta_{n},d_{n}),$ (9)
$\{\bm{q}_{n,1},...,\bm{q}_{n,K}\}=F_{q}(f_{n},\delta_{n},d_{n}),$ (10)
where the output of $F_{s}$ and $F_{q}$ are activated by sigmoid and normalization respectively. Furthermore, the overall geometric configuration of each cell is explicitly associated with that of its constituent Gaussians through
$\bm{s}_{n,k}=\mathcal{O}_{n,k}^{\text{scale}}*\bm{S}_{n},$ (11)
$\bm{Q}_{n}=\text{norm}(\sum_{k}^{K}\bm{q}_{n,k}*\sigma_{n,k}),$ (12)
To restrict that the Gaussian regions fully locate inside the cell region, the maximum of $\mathcal{O}_{n,k}^{\text{scale}}$ is clamped by
[…]
As illustrated in Fig. 3, the proposed neural Gaussians are constrained to reside strictly within their corresponding cell regions. Leveraging this property, the cell projections can effectively guide the Gaussian blending process.
IV-B Cell Searching Rasterization
In the original 3DGS framework [7], views are rendered by rasterizing 3D Gaussians onto the image plane through a Gaussian search rasterization strategy. In contrast, the anchors introduced in [16], which form the proxies of the dual-layered hierarchy, do not directly participate in the rendering process; instead, rendering remains confined strictly to the Gaussian level. This results in a distribution mismatch between the two hierarchy levels, i.e., between anchors and 3D Gaussians.
Building on our proposed geometric cell, we introduce a cell search rasterization technique in which cells, rather than individual Gaussians, serve as the primary rasterization units. As illustrated in Fig. 3, we perform rasterization at the cell level during the pre-processing stage, which reduces memory access overhead. Specifically, each cell defines a visible region, and only those whose regions intersect with the viewing frustum are recorded and subsequently decoded into neural 3D Gaussians for the subsequent $\alpha$-blending stage. This strategy ensures that the distribution of cells and 3D Gaussians are naturally aligned across the two hierarchy levels, and restricts rasterization exclusively to the cell level. Following the tile based paradigm, cells are assigned to individual tiles, and radix sort is applied to the cell sequence as necessary. Because each cell encapsulates a local cluster of Gaussians, the total number of cells is substantially less than the number of individual Gaussians. As a result, both memory consumption and computational overhead during rasterization are significantly reduced.
Figure 4: Comparison of three different transmittance: LC-WSR (Eq. 8), vanilla 3DGS (Eq. 3) and ours (Eq. 16). The proposed transmittance is not strictly monotonically decreasing with increasing depth, indicating that the Gaussians are locally unordered. Nevertheless, our method achieves blending weights similar to those of vanilla 3DGS, without requiring an explicit Gaussian order.
V Layer II: Physical Weighted Sum Rendering
Duplex-GS reconstructs 3D scenes using ellipsoidal cells, enabling rasterization and sorting to be performed at the cell level, thereby significantly reducing rasterization overhead. However, without explicitly sorting the depths of individual Gaussians—a requirement in the original $\alpha$-blending operations grounded in physical modeling—the exact blending order becomes ambiguous. Sort-free Gaussians [12], an OIT extension of vanilla 3DGS, demonstrated that optimized components and parameters for high-fidelity view synthesis are governed solely by the mathematical constraints of the scene model. This weighted sum approach is well-suited for integrating 3DGS into modern GPU graphics pipelines and for deployment on resource-constrained devices. Nevertheless, abandoning physical constraints introduces “transparency” artifacts that degrade visual quality, and the absence of early termination further results in reduced rendering efficiency.
In this section, we integrate the OIT based rendering paradigm with a cell based Gaussian model, retrieving physical constraints for WSR based 3DGS.
V-A Physical OIT Blending for 3DGS
WSR [12] is a mathematically motivated rendering model, in which $\alpha_{i}$ in Eq. 6 does not directly represent the meaningful opacity of a Gaussian and may even exceed one due to its lack of a physical basis. While this formulation facilitates the weighted sum rendering process, it inherently precludes the early-termination mechanisms used in conventional $\alpha$-blending.
Obtaining the sorted order at the cell level is both memory and computationally efficient, since the number of cells is substantially fewer than that of the 3D Gaussians. Leveraging this, we redefine $\alpha$ as the opacity weighted by the Gaussian kernel, as specified in Eq. 2, constraining its value to the physically meaningful range of $[0,1]$.
Thanks to the approximate ordering provided by cell-level sorting—which adequately reflects the relative relationships among the contained Gaussians—we are able to enforce physical constraints within the mathematical WSR framework, without the need for precise ordering of all Gaussians. As illustrated in Fig. 3, the final rendered view $\bm{C}$ is computed according to the weighted sum rendering formula:
$\bm{C}=\frac{\sum_{n=1}^{N}w_{n}\sum_{k=1}^{K}\alpha_{n,k}\bm{c}_{n,k}}{\sum_{n=1}^{N}w_{n}\sum_{k=1}^{K}\alpha_{n,k}},$ (14)
where $w_{n}$ denotes the weight of cell-$n$, derived from the cell-level sorting results. A sophisticated WSR kernel for each cell, $w_{n}$ can then be defined as:
$w_{n}=v_{n}\cdot T_{n}^{\text{cell}},$ (15)
where $T_{n}^{\text{cell}}$ denotes the cell-level transmittance, which is efficiently obtained by sorting the sparsely distributed cells and can be calculated as:
$T_{n}^{\text{cell}}=\prod_{i=1}^{n-1}\prod_{k=1}^{K}(1-v_{i}\cdot\alpha_{i,k}).$ (16)
Due to the non-uniform distribution and potential overlap of cells, $T_{i}^{\text{3DGS}}$ in Eq. 3 may not precisely capture the transmittance of all emitted neural Gaussians. The introduction of the learnable parameter $v_{n}$ for each cell effectively compensates for this approximation. Three relevant paradigms of transmittance and the resulting Gaussians weights are visualized in Fig. 4. LC-WSR simulates the decay trend of transmittance with a linear function w.r.t. Gaussian depth, introducing great number of Gaussians with tiny weights. Our method obtain similar weights compared with the vanilla 3DGS without the requirements of concrete order of individual Gaussians.
V-B Early Termination for Cell Based 3DGS
We propose to perform sorting based on cells, rather than individual Gaussians, thereby significantly reducing sorting overhead while retaining the hardware-friendly weighted sum rendering paradigm.
As $T_{n}^{\text{cell}}$ decreases during the front-to-back rendering of cells, early termination can be efficiently incorporated to accelerate the rendering process and mitigate transparency artifacts. Specifically, $T_{n}^{\text{cell}}$, defined in Eq. 16, approaches zero as cell-level transmittance accumulates and saturates, which provides a practical criterion for early termination: the blending process is halted when $T_{n}^{\text{cell}}<\epsilon$. Furthermore, in contrast to previous WSR methods for 3DGS, our approach restores explicit physical meaning to the opacity parameter $\sigma_{n,k}$, restricting it to the range $[0,1]$. This property facilitates pruning and regularization during training.
VI Optimization Strategies for Cell Based Gaussian Splatting
The proposed cell primitives are endowed with distinct physical attributes; however, these features are not directly involved in the differentiable blending process. Furthermore, advanced optimization strategies are introduced to systematically organize the evolving cell structures.
VI-A Alignment of Cell and Gaussian Region
The cell search rasterization inevitably leads to the introduction of redundant 3D Gaussians during the blending stage, due to the discrepancies that arise between the spatial intervals of cells and the emitted neural Gaussians. Furthermore, the efficiency of cell search rasterization is contingent upon the degree of coherence between the visible regions of a cell and its associated Gaussians.
VI-A1 False positive skip
Figure 5: Illustration of mismatch between cell and Gaussian. (a) The interval between layers influence the efficiency of cell rasterization. (b) False positive regions are skipped to alleviate the speed degradation from the mismatch.
TBR is widely adopted in modern GPUs and is also utilized in 3DGS rasterization. However, this approach often introduces a substantial number of redundant false-positive Gaussians, a problem that is further exacerbated in cell based Gaussian frameworks due to the spatial mismatch between cell regions and the actual coverage of Gaussians. To mitigate the associated performance degradation, we introduce a pixel-level radius based culling strategy to eliminate unnecessary Gaussian kernel computations. Specifically, as illustrated in Fig. 5, a lightweight axis-aligned bounding box (AABB) check is employed: Gaussians are excluded from the blending operation if the pixel lies outside their respective AABBs. This efficient early-exit mechanism significantly reduces redundant computation and accelerates the overall rendering process.
VI-A2 Geometric correction via neural Gaussians
In addition to skipping false-positive Gaussians to improve efficiency, maximizing the overlap between each cell and its emitted neural Gaussians further reduces the number of redundant Gaussians. To achieve better geometric alignment, we dynamically define the cell’s quaternion as a weighted average of the quaternions associated with its emitted Gaussians, as described in Eq. 12. Moreover, to ensure an even spatial distribution of Gaussians within each cell, the cell centers are periodically updated according to
$\bm{x}_{n}=\frac{\sum_{k=1}^{K}\bm{\mu}_{n,k}}{K}.$ (17)
As cell scales are differentiable via Eq. 11, the three-dimensional covariances and positions of first-layer cells are learned in accordance with the updates to the second-layer Gaussians.
VI-A3 Cell position reset
As discussed in Gaussian based models [7, 16, 17], the initial SfM point cloud often fails to comprehensively cover the entire scene, particularly in textureless or sparsely observed regions. Following the adaptive density control (ADC) policies employed in previous works [7], we utilize view-space positional gradients as a densification score to guide the growth of new cells. Specifically, additional cells are introduced at the centers of Gaussians that attain high densification scores. In contrast to the anchors proposed in [16, 17], which do not participate directly in the rendering process, our cells possess explicit physical attributes and independently encode local regions. To decouple existing cells from newly introduced ones, the offsets of the significant Gaussians are reset to zero after cell growth, i.e., their positions are reinitialized to the centers of the corresponding cells.
VI-B Perceptual Loss
The loss function for optimizing 3D Gaussian Splatting (3DGS) is typically formulated as a combination of $\ell_{1}$ and SSIM [33] losses, which respectively capture pixel-level reconstruction error and structural similarity. Minimizing these terms promotes high signal-to-noise ratios but often results in images that lack high-frequency details and appear perceptually unsatisfying, particularly at higher resolutions. Given the effectiveness of perceptual loss functions in image generative models [34, 35], we additionally incorporate a perceptual loss term for training Gaussian based models. The overall loss is defined as
$\mathcal{L}=(1-\lambda_{1}-\lambda_{2})\mathcal{L}_{1}+\lambda_{1}\mathcal{L}_{\text{SSIM}}+\lambda_{2}\mathcal{L}_{\text{LPIPS}},$ (18)
where $\mathcal{L}_{1}$, $\mathcal{L}_{\text{SSIM}}$ and $\mathcal{L}_{\text{LPIPS}}$ denote $\ell_{1}$, structural similarity and perceptual loss respectively. While the inclusion of the perceptual loss yields rendering results with improved perceptual quality, its computation relies on convolutional neural networks (CNNs), which increases the training time. To address this, we compute the perceptual loss every 20 training iterations, which substantially alleviates the additional computational overhead without noticeably increasing the overall training time.
VII Experimental Results
TABLE I: Quantitative comparisons on real-world dataset [36, 37, 38]. Duplex-GS achieves competitive rendering quality while reducing model size compared to the baselines. The best performance for each metric is highlighted.
Method Mip-NeRF360 Tanks & Temples Deep Blending
$\text{PSNR}_{\uparrow}$ $\text{SSIM}_{\uparrow}$ $\text{LPIPS}_{\downarrow}$ $\text{Storage}_{\downarrow}$ $\text{PSNR}_{\uparrow}$ $\text{SSIM}_{\uparrow}$ $\text{LPIPS}_{\downarrow}$ $\text{Storage}_{\downarrow}$ $\text{PSNR}_{\uparrow}$ $\text{SSIM}_{\uparrow}$ $\text{LPIPS}_{\downarrow}$ $\text{Storage}_{\downarrow}$
Mip-NeRF3601 [36] 27.69 0.792 0.237 - 22.22 0.759 0.257 - 29.40 0.901 0.245 -
3DGS1 [7] 27.52 0.813 0.222 750.2 MB 23.57 0.845 0.180 431.4 MB 29.61 0.900 0.251 662.7 MB
Mip-Splatting1 [39] 27.61 0.816 0.215 838.4 MB 23.96 0.856 0.171 500.4 MB 29.56 0.901 0.243 736.8 MB
[…]
LC-WSR2 [12] 27.21 0.800 0.219 1312.0 MB 23.22 0.832 0.186 672.7 MB 29.90 0.901 0.243 810.7 MB
Ours $(K=10)$ 27.90 0.813 0.216 152.2 MB 24.26 0.867 0.150 124.8 MB 30.30 0.910 0.249 89.7 MB
Ours $(K=5)$ 27.74 0.802 0.218 104.4 MB 24.32 0.867 0.138 112.2 MB 30.26 0.909 0.236 94.7 MB
  • 1
    Experiments are conducted using the official public repository, with no modifications except for iteration settings as described in Sec. VII-A4.
  • 2
    The official implementation has not been released. Experiments are conducted using our own implementation, which is available at https://github.com/LiYukeee/sort-free-gs.
TABLE II: Quantitative comparisons on BungeeNeRF [40] and VR-NeRF [1] datasets. Additional experiments assess performance on multi-scale rendering and the reconstruction of intricate indoor details. Duplex-GS achieves high-quality rendering with compact model sizes. The best result for each metric is highlighted.
Method BungeeNeRF VR-NeRF: apartment VR-NeRF: kitchen
$\text{PSNR}_{\uparrow}$ $\text{SSIM}_{\uparrow}$ $\text{LPIPS}_{\downarrow}$ $\text{Storage}_{\downarrow}$ $\text{PSNR}_{\uparrow}$ $\text{SSIM}_{\uparrow}$ $\text{LPIPS}_{\downarrow}$ $\text{Storage}_{\downarrow}$ $\text{PSNR}_{\uparrow}$ $\text{SSIM}_{\uparrow}$ $\text{LPIPS}_{\downarrow}$ $\text{Storage}_{\downarrow}$
3DGS1 [7] 27.71 0.915 0.099 1654.1 MB 30.98 0.922 0.212 368.1 MB 31.73 0.933 0.185 443.6 MB
Mip-Splatting1 [39] 28.21 0.922 0.096 1108.8 MB 31.49 0.929 0.201 442.8 MB 32.19 0.939 0.177 484.3 MB
[…]
LC-WSR2 [12] 26.94 0.903 0.111 1670.2 MB 31.84 0.935 0.169 1543.2 MB 32.32 0.937 0.158 1482.7 MB
Ours $(K=10)$ 28.80 0.920 0.093 319.2 MB 31.79 0.927 0.218 191.7 MB 32.17 0.934 0.191 176.2 MB
Ours $(K=5)$ 28.66 0.914 0.098 262.8 MB 32.27 0.932 0.178 299.0 MB 32.38 0.940 0.157 269.8 MB
  • 1
    Experiments are conducted using the official public repository, with no modifications except for iteration settings as described in Sec. VII-A4.
  • 2
    The official implementation has not been released. Experiments are conducted using our own implementation, which is available at https://github.com/LiYukeee/sort-free-gs.
VII-A Experimental Settings
VII-A1 Datasets
We conducted extensive experiments on 31 realistic scenes drawn from various datasets, following the configurations in 3DGS [7] and Scafolld-GS [16]. Specifically, we evaluated 13 publicly released datasets, including nine from Mip-NeRF360 [36], two from Tanks and Temples [37] and two from DeepBlending [38]. Additionally, we tested 10 further scenes—eight from BungeeNeRF [40] and two from VR-NeRF [1]—to assess the robustness of our approach in relatively large-scale environments. Scenes from BungeeNeRF feature multi-scale outdoor city scenes for validating view-adaptive rendering, while those from VR-NeRF provide intricate indoor scenes to evaluate detailed rendering capabilities.
VII-A2 Metrics
We evaluate the reconstruction fidelity with PSNR, SSIM [33] and LPIPS [41], which are commonly used in image quality measurement. We also record the model size, frame per second (FPS) as indicators of storage overhead and rendering speed. Additionally, the consumptions of the radix sort are listed to address the effectiveness of our method in reducing the sorting overhead, which is a bottleneck for edge devices.
VII-A3 Baselines
We compare our method with several SOTA approaches, including 3DGS, Mip-NeRF360, Scaffold-GS, Octree-GS, and LC-WSR. Specifically, LC-WSR serves as the baseline for OIT based Gaussian Splatting rendering, while the remaining methods employ the $\alpha$-blending paradigm used in the vanilla 3DGS.
VII-A4 Implementation Details
For the baseline models, we utilize the sum of the $\ell_{1}$ and SSIM losses as the loss function, with the weight $\lambda_{1}$ in Eq. 18 is set to $0.2$. For comparison, the perceptual loss is computed every 20 training iterations, with $\lambda_{2}$ set to $0.5$. For neural Gaussian based models (Scaffold-GS, Octree-GS, and the proposed Duplex-GS), experiments are conducted with encoding dimensions $K$ set to $5$ and $10$ respectively.
For small-scale scenes containing hundreds of training images, training is performed for 40,000 iterations, with densification ending after 25,000 iterations. For large-scale scenes containing thousands of training images, training extends to 100,000 iterations, and densification concludes at 50,000 iterations. All experiments are conducted on a single NVIDIA RTX 4090 GPU.
VII-B Results Analysis
We first validate our methods by showing that it achieves the highest accuracy across several datasets. The results are shown in Tab. I and Tab. II.
Figure 6: Qualitative comparisons of Duplex-GS and competing baselines on Mip-NeRF360 [36] dataset. The proposed Duplex-GS successfully removes the “transparency” artifacts for the WSR rendering kernel while maintaining the ability to reconstruct fine details.
VII-B1 Quantitative Results
Tab. I presents a comprehensive comparison of our method with several SOTA baselines on three real-world datasets. On the Mip-NeRF dataset, our approach achieves the highest PSNR when $K=10$, outperforming all baselines. While Mip-Splatting and Octree-GS attain slightly higher SSIM values (0.816), our method yields competitive SSIM (0.813) and achieves improved perceptual quality, as evidenced by a lower LPIPS (0.216). Remarkably, our model also demonstrates strong storage efficiency, requiring significantly less memory than methods such as 3DGS and Mip-Splatting, both of which demand over 750 MB.
On the Tanks&Temples and Deep Blending datasets, although the PSNR values of our approach are marginally lower, it delivers superior visual quality, attaining the highest SSIM and markedly lower LPIPS compared to other models. Notably, when compared to LC-WSR—the OIT variant of Gaussian Splatting—our method achieves better performance across all metrics and datasets, showing the effectiveness of the OIT paradigm in Gaussian Splatting on CUDA based high-end GPUs.
To further assess our method under challenging scenarios involving multi-scale training views and large-scale datasets—common in real-world applications—we conduct experiments on the BungeeNeRF and VR-NeRF datasets (results shown in Table II). On BungeeNeRF, our approach achieves the highest PSNR and lowest LPIPS when $K=10$, and remains highly competitive with $K=5$, attaining the smallest model size among all baselines. The two scenes within VR-NeRF are large-scale, intricate indoor environments containing thousands of training images. Experimental results indicate that OIT based rendering consistently outperforms sort based $\alpha$-blending approaches for such cases, and our proposed method achieves the highest overall accuracy among all evaluated baselines.
VII-B2 Qualitative Results
We present qualitative results and comparisons with SOTA baselines to demonstrate the effectiveness of our proposed method in artifact removal and detail preservation. As illustrated in Fig. 6 using examples from the Mip-NeRF360 dataset, our method eliminates the “transparency” artifacts observed in the previous OIT based Gaussian model, LC-WSR, while maintaining intricate scene details. For multi-scale and large-scale scenarios, as shown in Fig. 7, the proposed Duplex-GS successfully captures fine details, even in textureless and marginal regions. Additionally, our approach effectively removes “popping” artifacts, benefiting from the OIT blending kernel (see Fig. 8).
Figure 7: Qualitative comparisons of Duplex-GS and competing baselines on BungeeNeRF [36] and VR-NeRF [1] datasets. the proposed Duplex-GS successfully captures fine details, even in textureless and marginal regions.
TABLE III: Efficiency analysis of Duplex-GS and competing baselines.
Method Mip-NeRF360 Tan.&Tem. Deep Blend. BungeeNeRF
3DGS [7] 149 142 145 52
Scaffold-GS [16] 150 129 185 72
Octree-GS [17] 183 159 271 114
LC-WSR [12] 77 89 114 31
Ours 184 147 232 124
VII-B3 Efficiency Analysis
Tab. III provides a comparison of rendering speeds, measured in FPS, between our method and several state-of-the-art baselines across four benchmark datasets. Our approach consistently achieves the highest or highly competitive FPS on all datasets. On the Mip-NeRF360 dataset, our method achieves 184 FPS, surpassing all other baselines and comparable to the fastest prior approach, Octree-GS (183 FPS). For the Tanks & Temples dataset, our method obtains 147 FPS, demonstrating performance close to the best-performing baseline. While on the Deep Blending dataset, our FPS of 232 is similarly on par with the leading method. Most notably, on the BungeeNeRF dataset, our method achieves 124 FPS, a considerable improvement over existing baselines. These results highlight the efficiency and scalability of our approach, demonstrating real-time rendering capability and practical superiority in intricate and multi-scale scenarios.
Radix sort is employed in 3DGS to determine the explicit ordering of Gaussian primitives, sorting $b$-digit numbers with $m$ possible values in $\mathcal{O}\left(b\cdot(n+m)\right)$ time and $\mathcal{O}(n+m)$ space. Notably, the sequence length $n$ contributes linearly to the overall complexity. To evaluate the effectiveness of our approach in reducing the computational and memory requirements of view-adaptive sorting, we record the sorting length and corresponding memory consumption during the radix sort procedure. The results, presented in Tab. IV, show that our method successfully mitigates the sorting overhead by introducing cell proxies, thereby demonstrating superior efficiency compared to existing approaches.
TABLE IV: Comparisons about the overhead of radix sort stage. The proposed cell rasterization consistently reduce both the time and space consumptions compared with the vanilla 3DGS rasterization.
Method Mip-NeRF360 BungeeNeRF
$\text{Sorting Length}_{\downarrow}$ $\text{Memory}_{\downarrow}$ $\text{Sorting Length}_{\downarrow}$ $\text{Memory}_{\downarrow}$
3DGS [7] 7,000,602 81.0 MB 21,798,876 255.0 MB
Ours 3,217,845 37.7 MB 2,843,604 33.3 MB
VII-C Ablation Studies
VII-C1 Early Termination
The introduction of the early termination mechanism is a key advantage of our method over previous approaches such as LC-WSR. To visualize the mechanism, we record the termination ratio with
$r^{\text{ET}}=1-\frac{N^{\text{rendered}}}{N^{\text{val}}},$ (19)
where $N^{\text{rendered}}$ denotes the number of Gaussians contributed to the final color, $N^{\text{val}}$ denotes that of the total interacted valid Gaussians along the ray. The results are visualized in Fig. 9. Our method demonstrates termination behavior which closely matches that of the vanilla 3DGS, whereas prior sort-free Gaussian Splatting methods lack this feature. To evaluate its effectiveness, we conduct an ablation study by disabling early termination during model training. As shown in Tab. V, enabling early termination leads to improvements in both accuracy and efficiency.
Figure 8: Illustration of “popping” artifacts removal by recording the differences between contiguous timestep.
Figure 9: Visualization of the early termination mechanism. The existing OIT 3DGS method, LC-WSR, does not support early termination due to its sort-free design. In contrast, Duplex-GS achieves termination results comparable to those of vanilla 3DGS, without requiring an explicit Gaussian order.
TABLE V: Ablation experiments.
Method $\text{PSNR}_{\uparrow}$ $\text{SSIM}_{\uparrow}$ $\text{LPIPS}_{\downarrow}$ $\text{FPS}_{\uparrow}$
w/o early termination 27.66 0.801 0.220 155
w/o cell restriction 27.54 0.799 0.223 229
w/o $v_{i}$ 27.59 0.801 0.218 174
Full method 27.74 0.802 0.218 184
VII-C2 Cell Geometry Restriction
Compared with anchor, cell has specific geometry, restricted by Eq. 12 and Eq. 13, to align the regions between cell and the corresponding Gaussians. To show the effects of these conditions, we mute these restriction and the results are shown in Tab. V. The accuracy decreases as expected, however, the rendering speed is improved, which is regarded as the compromise between accuracy and efficiency.
VII-C3 Transmittance Design
As the sorting results of cell level cannot reflect the exact order of Gaussians. we introduce learnable parameter $v_{i}$ to compensate the ambiguity. We train models without $v_{i}$ to show its effectiveness. The results are shown in Tab. V. By introducing $v_{i}$, the accuracy is improved and the efficiency is slightly improved due to the different amount of cell primitives.
VIII Limitation and Conclusion
In this work, we propose a dual-layered Gaussian model that effectively resolves the “popping” artifacts present in vanilla 3DGS and the “transparency” artifacts associated with OIT Gaussian Splatting. Leveraging cell search rasterization and an OIT weighted blending technique, our approach substantially reduces the overhead of global sorting. Implemented with CUDA, our method reconstructs intricate scene details while maintaining real-time rendering speed, thereby demonstrating the practicality and effectiveness of the OIT rendering paradigm for Gaussian Splatting models.
Nevertheless, several limitations remain. Specifically, the geometric relationships between cells and Gaussians are not yet fully explored, leading to performance degradation due to intervals as illustrated in Fig. 5. Additionally, while our current implementation is based on CUDA devices, it does not yet exploit intra-cell parallelism for Gaussian blending. Future work will aim to address these limitations, including efficient implementation on specialized hardware such as Field-Programmable Gate Array (FPGA) and Application-Specific Integrated Circuit (ASIC).

### v2

@@
 Abstract
-Recent advances in 3D Gaussian Splatting (3DGS) have demonstrated remarkable rendering fidelity and efficiency.
-However, these methods still rely on computationally expensive sequential $\alpha$-blending operations, resulting in significant overhead—particularly on resource-constrained platforms.
-In this paper, we propose Duplex-GS, a dual-hierarchy framework that integrates proxy Gaussian representations with order-independent rendering techniques to achieve photorealistic results while sustaining real-time performance.
-To mitigate the overhead caused by view-adaptive radix sort, we introduce cell proxies for local Gaussians management and propose cell search rasterization for further acceleration.
-By seamlessly combining our framework with Order-Independent Transparency (OIT), we develop a physically inspired weighted sum rendering technique that simultaneously eliminates “popping” and “transparency” artifacts, yielding substantial improvements in both accuracy and efficiency.
-Extensive experiments on a variety of real-world datasets demonstrate the robustness of our method across diverse scenarios, including multi-scale training views and large-scale environments.
-Our results validate the advantages of the OIT rendering paradigm in Gaussian Splatting, achieving high-quality rendering with an impressive $1.5\times$$4\times$ speedup over existing OIT based Gaussian Splatting approaches and $52.2\%\,$$86.9\%$ reduction of the radix sort overhead without quality degradation.
-Index Terms:
-Novel View Synthesis, Gaussian Splatting, Order-Independent Transparency, Real-Time Rendering
+3D Gaussian Splatting (3DGS) achieves photorealistic rendering but incurs substantial overhead due to the global sorting required for $\alpha$-blending, which results in noticeable “popping” artifacts and hinders deployment on edge devices.
+While sort-free Order-Independent Transparency (OIT) methods circumvent sorting, they introduce “transparency” artifacts and suffer from inefficiency due to the absence of physical constraints.
+To address these limitations, we present Duplex-GS, a dual-hierarchy framework that leverages proxy-guided spatial organization and a novel hybrid renderer that combines $\alpha$-blending with reformulated Weighted-Sum Rendering (WSR).
+We introduce explicit ellipsoidal cell proxies to encapsulate local Gaussians, which enables efficient proxy-level rasterization.
+This strategy drastically reduces the overhead associated with global sorting.
+Furthermore, we propose a physically grounded WSR scheme with cell-level early termination, which retrieves the physical constraints absent in prior OIT-based 3DGS methods, effectively eliminating both popping and transparency artifacts.
+Extensive experiments on diverse real-world datasets demonstrate the robustness of Duplex-GS across scenarios spanning multi-scale training views and large-scale environments.
+Quantitatively, our method delivers high-fidelity real-time rendering, outperforming prior OIT-based 3DGS methods by $1.5\times$– $4\times$ in speed, while reducing radix-sort cost by $29.8\%\,$– $86.9\%$ compared with conventional $\alpha$-blending without compromising visual quality.
+publicationid: pubid: 0000–0000/00\$00.00 © 2021 IEEE
 I Introduction
-Novel view synthesis (NVS) and real-time rendering have attracted significant attention in both academia and industry, with broad applications in video games, virtual reality (VR) [1], and 3D scene reconstruction [2, 3, 4, 5].
-Inspired by the groundbreaking Neural Radiance Field (NeRF) [6], learning based methods have substantially advanced the realism of digital content, narrowing the gap between the virtual and real worlds.
+Novel view synthesis (NVS) and real-time rendering have attracted significant attention in both academia and industry, with broad applications in video games, virtual reality (VR) [1], and 3D scene reconstruction [2, 3, 4, 5, 6, 7, 8, 9].
+Inspired by the groundbreaking Neural Radiance Field (NeRF) [10], learning-based methods have substantially advanced the realism of digital content, narrowing the gap between the virtual and real worlds.
 However, these methods often entail considerable computational overhead.
-The emergence of 3D Gaussian Splatting (3DGS) [7, 8] marks another milestone, enabling photorealistic and real-time rendering of radiance fields via rasterization—a paradigm akin to the graphics pipeline in modern graphics processing units (GPUs).
-Despite the impressive reconstruction and rendering capabilities of 3DGS, a primary limitation is the need for an ever-increasing number of Gaussian primitives to fit all training views.
+The emergence of 3D Gaussian Splatting (3DGS) [11, 12] marks another milestone, enabling photorealistic and real-time rendering of radiance fields via rasterization, a paradigm akin to the graphics pipeline in modern graphics processing units (GPUs).
+Despite the impressive reconstruction and rendering capabilities of 3DGS, a primary limitation is its need for an ever-increasing number of Gaussian primitives to fit all training views.
 This results in substantial and irregular memory access during rendering, as costly view-dependent sorting operations over long sequences impose considerable computational and memory overhead [13, 14].
@@
 More recently, sort-free Gaussian Splatting [16] has incorporated Order-Independent Transparency (OIT) [17, 18] into the Gaussian Splatting framework, eliminating the need for sorting in the rasterization pipeline by employing Weighted Sum Rendering (WSR) [19].
-This effectively eliminates “popping” artifacts and has been successfully demonstrated on mobile devices [12].
+This effectively eliminates popping artifacts and has been successfully demonstrated on mobile devices [16].
 However, as illustrated in Fig.
-1, the absence of early termination when encountering opaque objects leads to “transparency” artifacts and increased rendering time, particularly in large-scale scenes, due to the accumulation of redundant Gaussian primitives in final blending.
-Neural Gaussian framework [16, 17] is another line of work to enhance storage efficiency for 3DGS, which transitions from explicit 3D Gaussian primitives to a hybrid representation where clusters of Gaussians are encoded via “anchors” and associated with a shared MLP.
-However, this paradigm adheres to the conventional rendering pipeline and still operates on sets of 3D Gaussians decoded on-the-fly by the MLP.
-As a result, sorting the lengthy sequence of Gaussians remains mandatory for such Tile based Rendering (TBR), presenting substantial challenges for devices with limited resources [10].
-Furthermore, since the spatial positions of neural Gaussians are not explicitly constrained by the anchor points, the generated Gaussians may deviate significantly from their associated anchors—as illustrated in Fig.
-2—thereby complicating anchor management.
+1, the absence of physical constraints leads to “transparency” artifacts and increased rendering time, particularly in large-scale scenes, due to the accumulation of redundant Gaussian primitives and lack of early termination when encountering opaque objects in the blending phase.
+Neural Gaussian frameworks [20, 21] enhance storage efficiency by using “anchor” points to encode local Gaussians via shared MLPs.
+However, they still require sorting the decoded Gaussians for $\alpha$-blending, which is costly on resource-constrained devices [14].
+Moreover, the lack of explicit spatial constraints allows Gaussians to deviate from their anchors (Fig.
+2), complicating the representation.
 Figure 1: Motivation of Duplex-GS.
-The rendering process for 3DGS rely on sorting for long tile-Gaussian sequence, introducing popping artifacts.
-The sort-free LC-WSR uses weighted sum rendering method without the need of consuming sorting stage, introducing transparency artifacts due to the lack of early termination when encounter opaque object.
+(Top) 3DGSs are rendered via $\alpha$-blending for a long tile-Gaussian sequence.
+However, sudden changes in viewing angle can cause abrupt changes in the blending order and the corresponding weights, resulting in noticeable popping artifacts.
+(Bottom) The sort-free LC-WSR method eliminates the sorting stage through WSR paradigm.
+However, the absence of physical constraints leads to transparency artifacts and efficiency degradation, as blending continues unnecessarily for opaque surfaces.
 Figure 2: Overview of the proposed Duplex-GS.
-The dual-hierarchy structure introduce cell as the proxy of local similar Gaussians.
-(a) Compared with proxy anchor, the proposed proxy cell is an ellipsoid with specific geometry.
-(b) Combing OIT rendering with 3DGS models, our method successfully remove the “popping” and “transparency” artifacts simultaneously.
-(c) By introducing cell search rasterization (Sec.
-IV) and physical weighted sum rendering (Sec.
-V), we achieve significant improvement over OIT Gaussian Splatting for both accuracy and efficiency.
-In this paper, we present Duplex-GS, which models 3D scenes with WSR rendering paradigm and dual-layered structure following the neural Gaussian framework.
-Firstly, scenes are encoded as explicit proxies that are easily stored and manipulated in the first layer.
-As the concept anchor degenerates the 3D Gaussian to point cloud, we introduce “cell” as the replacement, which is an ellipsoid with concrete physical geometry initialized at the Structure-from-Motion (SfM) [18] points.
-By constraining the emitted neural Gaussians located inside the cell region, the rasterization can be performed in cell level.
-Secondly, primitives are decoded as other forms, which is suitable for high-fidelity rendering in the second layer.
-Encouraged by [12], we introduce the physical inspired WSR with early termination where the specific order of Gaussian depth is not required, which replaces the non-commutative blending operations with parallelizable weighted sum operation.
-As the number of primitives in the first layer is extensively fewer than that of Gaussians, the proposed cell search rasterization significantly reduce the memory and computational overhead introduced by radix sort.
-On the one hand, rasterization and lightweight sorting in the first layer removes the “transparency” artifact.
-On the other hand, the WSR kernel removes the “popping” artifact introduce by vanilla $\alpha$-blending.
-In summary, the contributions of this work are as follow:
-  • We introduce a dual-layered Gaussian hierarchy, in which the first layer consists of proxies called cells and the second layer comprises 3D Gaussians.
-Our approach significantly alleviates the need for global sorting while preserving the parallelizability of WSR, thereby achieving artifact-free, real-time 3D reconstruction.
-  • Cell rasterization in layer I: We introduce cells—ellipsoids with adaptive geometric properties—to organize emitted Gaussians.
-Based on this structure, we propose a novel cell rasterization strategy that dramatically reduces both memory and computational overhead associated with radix sort.
-  • Physical inspired WSR blending in layer II: We retrieve physical constraints of Gaussian parameters for WSR method to blend Gaussians without the need for global sorting, inherently avoiding the “popping” artifact.
-Building on this, we introduce cell-level early termination to further accelerate rendering and effectively mitigate the “transparency” artifact encountered in the original WSR approach.
-  • We implement WSR based Gaussian Splatting on CUDA-enabled platforms.
-Experimental results demonstrate that our method delivers comparable rendering speed and visual fidelity to existing approaches on commercial high-end GPUs.
+Our approach is built on a dual-hierarchy structure that introduces ellipsoidal cells as proxies for local Gaussians.
+(a) Explicit Cell Proxies.
+Unlike abstract anchor points, our geometrically defined proxy cells are directly utilized for rasterization, guiding the subsequent Gaussian blending process.
+(b) Hybrid Rendering Paradigm.
+By synergistically combining $\alpha$-blending with physically grounded WSR, our method simultaneously eliminates both popping and transparency artifacts.
+(c) The integration of cell rasterization (Sec.
+IV) and physically grounded WSR (Sec.
+V) achieves significant improvements over prior OIT-based Gaussian splatting methods in both rendering accuracy and computational efficiency.
+This paper presents Duplex-GS, a novel framework for modeling 3D scenes using explicit proxies with defined visible regions and a tailored hybrid rendering paradigm.
+Our approach integrates the benefits of $\alpha$-blending and WSR to eliminate artifacts while maintaining real-time performance.
+First, we encode scenes using ellipsoidal proxies, termed “cells”, which replace the anchors used in the original neural Gaussian framework.
+By constraining neural Gaussians to reside within their respective cells, rasterization is performed at the cell level to guide subsequent Gaussian blending.
+Second, activated cells are decoded into Gaussians via shared MLPs.
+By restoring physical constraints absent in previous OIT-based 3DGS [16], we introduce a physically grounded WSR strategy that enables early termination and replaces non-commutative blending operations with a parallelizable weighted sum.
+This hybrid design offers two key advantages.
+First, the significantly smaller population of cell proxies compared to Gaussians enables our cell-level rasterization strategy to drastically reduce the memory and computational overhead of global radix sorting.
+Second, we restore physical constraints absent in prior OIT-based 3DGS method, which facilitates cell-level early termination and removes both popping and transparency artifacts simultaneously.
+In summary, the contributions of this work are as follows:
+  • We propose a dual-layered Gaussian hierarchy comprising ellipsoidal proxy cells and 3D Gaussians.
+This structure enables artifact-free, real-time 3D reconstruction while maintaining the parallelizability of WSR.
+  • Cell rasterization (layer I): We introduce ellipsoidal cells with adaptive geometric properties to spatially organize neural Gaussians.
+A novel cell-level rasterization strategy is proposed, which significantly reduces the memory and computational overhead associated with global radix sorting.
+  • Physically grounded WSR blending (layer II): We incorporate physical constraints into the WSR method to blend local Gaussians, inherently avoiding popping and transparency artifacts.
+Building on this, we introduce a cell-level early termination scheme that further accelerates rendering.
+  • We provide a high-performance CUDA implementation of our WSR-based Gaussian splatting framework.
+Experimental results demonstrate that our method achieves competitive rendering speed and superior visual fidelity on consumer-grade hardware.
 II Related Work
 II-A Neural Representation for 3D Scene
-Traditional scene reconstruction methods, such as SfM [18, 19] and Multi-View Stereo (MVS) [20, 21], often struggle to recover texture-less regions or handle challenging lighting variations.
-The advent of Neural Radiance Fields (NeRF) [6] has demonstrated the impressive capability of learning based models to achieve photorealistic reconstructions from sparse 2D images.
-Building on this, numerous efficient NeRF variants have been developed [22, 23, 24], making real-time rendering increasingly feasible on edge devices [25].
-A significant breakthrough in the field is 3D Gaussian Splatting (3DGS) [7], which leverages explicit anisotropic 3D Gaussians for scene reconstruction and substantially accelerates rendering.
-Nevertheless, the substantial computational and memory demands of 3DGS continue to limit its applicability on resource-constrained platforms.
-Notably, a trend is witnessed that as scene representations become more explicit, rendering speed improves at the cost of increased memory requirements.
-And hybrid variants have emerged as a promising compromise between rendering efficiency and memory usage [22, 26, 16].
-For instance, Instant-NGP [22] employs an explicit learnable hash table that maps 3D space to a multi-resolution grid, thereby alleviating the computational burden on the MLP.
-Scaffold-GS [16] introduces a hierarchical structure that encodes neural Gaussians using anchor points, achieving comparable rendering quality and speed with a significantly reduced model size.
-These advances enable fast, high-resolution rendering without compromising visual fidelity.
-However, these approaches still rely on the conventional blending pipeline requiring the Gaussians to be rendered in sorted depth order, typically achieved with tile based sorting algorithms.
-This sorting step not only complicates practical implementation but also introduces visual artifacts such as sudden changes in appearance, commonly referred to as “popping” artifacts [11].
+Traditional scene reconstruction methods like SfM [22, 23] and MVS [24, 25] struggle with texture-less regions and lighting variations.
+NeRF [10] enabled photorealistic reconstruction from sparse images, with subsequent efficient variants [26, 27, 28] making real-time rendering feasible on edge devices [29, 30].
+3DGS [11] marked a breakthrough by using explicit anisotropic Gaussians for fast rendering, though its computational and memory demands limit its deployment on resource-constrained platforms.
+A clear trend shows that more explicit representations improve speed at the cost of memory, leading to hybrid approaches [26, 31, 32, 20, 33, 34, 35] that balance this trade-off.
+For example, Instant-NGP [26] uses a multi-resolution hash table to reduce MLP burden, Scaffold-GS [20] encodes neural Gaussians via anchors to reduce model size while maintaining quality.
+However, these methods still rely on the conventional $\alpha$-blending pipeline that requires sorting Gaussians by depth.
+This step complicates implementation and introduces visual popping artifacts [15].
 II-B Efficient Neural Rendering
-Unlike NeRF whose constraint is the high computational demand, the bottlenecks of 3DGS practical deployment are the extensive number of Gaussian primitives and the frequent memory access requirements [10, 9].
-Various works aim to improve the efficiency of 3DGS by reducing the model size or the number of Gaussians involved in the sorting based blending operations [17, 27, 28, 29].
-Encouraged by the OIT [13, 14], a widely deployed rendering paradigm in modern GPU to blend non-opaque media, sort-free Gaussian Splatting [12] introduce a novel blending paradigm WSR which eliminates the need of sorting operation in 3DGS, which removes the “popping” artifact as well as achieves faster rendering speed on mobile devices with small scale scenes.
-However, the lack of early termination results in the “transparency” artifacts as well as introduces extra computation when encounter opaque objects.
+While NeRF is primarily constrained by high computational demands, the practical deployment of 3DGS is bottlenecked by its extensive number of Gaussian primitives and frequent memory access requirements [14, 13].
+Existing efforts to optimize the original $\alpha$-blending pipeline in 3DGS broadly fall into three categories: 1) Applying model compression techniques, such as Level-of-Detail (LOD) strategy [21, 36, 37, 38], encoding schemes [20, 39, 40], as well as pruning and quantization [41, 42, 43].
+2) Optimizing critical operators through CUDA-based implementations [44, 45, 46, 47, 48] or dedicated hardware acceleration units [49].
+3) Replacing Gaussian kernels with more efficient representation such as linear kernals [50] and half-Gaussian kernel [51].
+A few works have attempted to revise the core rendering paradigm itself by exploring efficient alternatives to $\alpha$-blending.
+StochasticSplats [52] introduces a stochastic transparency paradigm which enables efficient rendering using OpenGL shaders at the risk of controllable quality degradation.
+Sort-free Gaussian Splatting [16] introduces a Weighted Sum Rendering (WSR) pipeline, which eliminates popping artifacts and achieves faster rendering on mobile devices for small-scale scenes.
+Nevertheless, its lack of physical constraints leads to transparency artifacts and introduces unnecessary computation when rendering opaque surfaces.
 II-C Order Independent Transparency
-Order-Independent Transparency (OIT) [13, 14] is a foundational graphics technique integrated into modern GPUs to enable the correct compositing of transparent fragments without requiring explicit depth sorting prior to rasterization.
+OIT [17, 18] is a foundational graphics technique integrated into modern GPUs to enable the correct compositing of transparent fragments without requiring explicit depth sorting prior to rasterization.
 Early approaches, such as Depth Peeling [53], A-buffer [54], and Multi-Layer Alpha Blending (MLAB) [55], successfully achieve OIT but often incur significant time or memory overhead.
 Weighted Blended OIT [19] offers a more efficient alternative by approximating the integration of transparency in a single render pass, storing weighted sums of color and opacity with carefully designed blending weights.
-And this method has been applied to 3D Gaussian Splatting (3DGS) rendering [12].
-However, the absence of early termination in [12] when encountering opaque regions can degrade both visual quality and rendering performance.
-In contrast, the proposed Duplex-GS framework integrates the core idea of early termination from depth peeling—halting compositing upon encountering fully opaque surfaces—with Weighted Blended OIT.
+And this method has been applied to 3DGS rendering [16].
+However, the absence of physical constraints in[16] degrades both visual quality and rendering efficiency.
+In contrast, the proposed Duplex-GS framework restores the physically grounded transmittance in proxy-level, enabling early termination to halt compositing upon encountering fully opaque surfaces with weighted blended OIT.
 As a result, Duplex-GS attains rendering quality and speed on par with state-of-the-art (SOTA) 3DGS models on high-end GPUs, while exhibiting strong potential for deployment on resource-constrained edge devices by alleviating the huge overhead of the view-adaptive sorting stage.
 Figure 3: Illustration of the proposed Duplex-GS pipeline.
-A dual-layered structure is designed to organize cells and Gaussians.
-In Layer I, proxy cells—represented as ellipsoids with specific geometric properties—are designed to aggregate Gaussians, enabling rasterization at the cell level and thereby alleviating the overhead associated with view-adaptive global sorting.
-Leveraging shared MLPs, the cell primitives are subsequently decoded into Gaussian primitives.
-In Layer II, the decoded Gaussians are blended using weighted sum rendering.
-Additionally, early termination technique is implemented—guided by the sorting results of the sparse cells from Layer I—which facilitates real-time, photorealistic rendering.
-III Preliminaries
-III-A Gaussian Splatting
-3D Gaussian Splatting [7] models scenes using anisotropic 3D Gaussian primitives and renders images through a rasterization based paradigm.
-Each 3D Gaussian $G(\bm{x})$ is represented by a center coordinate $\bm{\mu}\in\mathbb{R}^{3}$ and a covariance matrix $\bm{\Sigma}\in\mathbb{R}^{3\times 3}$:
-$G(\bm{x})=e^{-\frac{1}{2}(\bm{x}-\bm{\mu})^{T}\bm{\Sigma}^{-1}(\bm{x}-\bm{\mu})},$ (1)
-where $\bm{x}\in\mathbb{R}^{3}$ is the queried position of the Gaussian kernel, and the covariance $\bm{\Sigma}$ is parameterized by scales $\bm{s}\in\mathbb{R}^{3}$ and quaternions $\bm{q}\in\mathbb{R}^{4}$.
-In line with volumetric rendering conventions, each Gaussian is further associated with an opacity $\sigma\in\mathbb{R}$ and a view-dependent color $\bm{c}\in\mathbb{R}^{3}$, which is commonly encoded using Spherical Harmonics (SH).
-Rendering proceeds via a tile based Gaussian rasterization technique, which employs efficient radix sort to enable $\alpha$-blending
-$\bm{C}(\bm{x}^{\prime})=\sum_{i\in N}T_{i}^{\text{3DGS}}\bm{c}_{i}\alpha_{i},\quad\alpha_{i}=\sigma_{i}G^{\prime}_{i}(\bm{x}^{\prime}),$ (2)
-where $\bm{x}^{\prime}$ denotes the 2D coordinates of the queried pixel, $G^{\prime}_{i}(\bm{x}^{\prime})$ is the projected 2D gaussian kernel, and $T_{i}^{\text{3DGS}}$ denotes the transmittance
-$T_{i}^{\text{3DGS}}=\prod_{j=1}^{i-1}(1-\alpha_{i}).$ (3)
-III-B Scaffold-GS
-Scaffold-GS introduces anchors as local encoding proxies, which emit neural Gaussians decoded from latent features to effectively represent local scene structure.
+(Layer I) Ellipsoidal cells act as proxies to spatially organize local Gaussians.
+This structure enables efficient cell-level rasterization, significantly reducing the overhead of view-adaptive global sorting.
+(Layer II) Intersected cells are decoded into Gaussians and blended with the proposed physically grounded WSR.
+The final color is computed by first determining cell-level weights via $\alpha$-blending, then aggregating the Gaussians inside each cell in an order-independent WSR manner.
+III Preliminaries and Overview
+III-A Preliminaries
+III-A1 Neural Gaussian
+3DGS [11] models scenes using anisotropic 3D Gaussian primitives and renders images through a rasterization based paradigm.
+Building upon 3DGS, Scaffold-GS [20] introduces anchors as local encoding proxies, which emit neural Gaussians decoded from latent features to effectively represent local scene structure.
 Specifically, each anchor emits a predefined set of $K$ neural gaussians, whose centers are given by:
@@
 $\{\sigma_{n,1},...,\sigma_{n,K}\}=F_{\sigma}(\bm{f}_{n},\delta_{n},\bm{r}_{n}),$ (2)
-where $F_{\sigma}$ is the shared opacity decoder.
-Similarly, the other attributes of each neural Gaussians, including color $\bm{c}_{n,k}$, scales $\bm{s}_{n,k}$ and quaternions $\bm{q}_{n,k}$, are predicted by the separate MLPs $F_{c}$, $F_{s}$, $F_{q}$, respectively.
-III-C Weighted Sum Rendering
-Inspired by Order Independent Transparency (OIT)—a widely adopted technique for rendering transparent media in traditional graphics pipelines—sort-free Gaussian Splatting integrates this concept into 3DGS, producing the final image as
+where $F_{\sigma}$ is the shared opacity decoder where the output is activated by Tanh.
+Similarly, the other attributes of each neural Gaussians, including color $\bm{c}_{n,k}$, scales $\bm{s}_{n,k}$ and quaternions $\bm{q}_{n,k}$, are predicted by separate MLPs $F_{c}$, $F_{s}$, $F_{q}$, respectively.
+Consequently, rendering proceeds via a tile based Gaussian rasterization technique, which employs radix sort to blend $M$ ordered points
+$\bm{C}(\bm{x}^{\prime})=\sum_{i\in M}T_{i}^{\text{$\alpha$-blending}}\bm{c}_{i}\alpha_{i},\quad\alpha_{i}=\sigma_{i}G^{\prime}_{i}(\bm{x}^{\prime}),$ (3)
+where $\bm{x}^{\prime}$ denotes the 2D coordinates of the queried pixel, $G^{\prime}_{i}(\bm{x}^{\prime})$ is the projected 2D gaussian kernel, and $T_{i}^{\text{$\alpha$-blending}}$ denotes the transmittance
+$T_{i}^{\text{$\alpha$-blending}}=\prod_{j=1}^{i-1}(1-\alpha_{j}).$ (4)
+III-A2 Weighted Sum Rendering
+Inspired by OIT, a widely adopted technique for rendering non-opaque media in traditional graphics pipelines, sort-free Gaussian Splatting integrates this concept into 3DGS, producing the final image as
 $\bm{C}=\frac{\bm{c}_{B}w_{B}+\sum_{i=1}^{N}\bm{c}_{i}\alpha_{i}w(d_{i})}{w_{B}+\sum_{i=1}^{N}\alpha_{i}w(d_{i})},$ (5)
 where $\bm{c}_{B}$ and $w_{B}$ denote the background color and the learnable background weights respectively, and $d_{i}$ indicates the depths of the $i$-th 3D Gaussians.
-By employing this WSR formulation, the non-commutative, computationally intensive sorting stage—traditionally required for compositing—is eliminated.
+By employing this WSR formulation, the non-commutative, computationally intensive sorting stage, traditionally required for compositing, is eliminated.
 This significantly enhances the potential for parallelization and accelerates rendering.
-Specifically, the SOTA Linear Correction weighted-sum rendering (LC-WSR) function $w(d_{i})$ is defined as
+Specifically, the Linear Correction Weighted-Sum Rendering (LC-WSR) function $w(d_{i})$ achieve best performance which is defined as
 $w(d_{i})=T_{i}^{\text{LC-WSR}}\cdot v_{i},$ (6)
@@
 where $\tau$ and $v_{i}$ are the learnable parameters. This function formulates the decay trend w.r.t. the depths of the rendering Gaussians.
-IV Layer I: Cell Rasterization for 3DGS
-While neural Gaussian frameworks [16, 17] have achieved remarkable reconstruction accuracy, the use of anchors as encoding proxies lacks a concrete geometric interpretation.
-So we redesign the neural Gaussian proxies to alleviate the computational overhead associated with sorting and rasterization.
-Specifically, instead of encoding 3D Gaussians with anchors, we introduce the concept of “cell”—an ellipsoid with explicit geometric properties and a defined visible region—to encode neural Gaussians together with their latent features.
-Based on this representation, we further propose a cell rasterization strategy, enabling efficient and lightweight rasterization.
+III-B Overview
+We introduce a novel framework featuring a tailored encoding and rendering paradigm, as illustrated in Fig.
+3.
+Our approach is structured in two primary stages.
+First, 3D scenes are modeled using ellipsoidal cells, which act as proxies that emit Gaussians with coherent features inside their coverage.
+This is coupled with a hybrid rendering paradigm that integrates the advantages of $\alpha$-blending and WSR, establishing a globally ordered yet locally unordered scene structure.
+Specifically, rasterization is first performed at the proxy layer to compute a blending weight for each cell in $\alpha$-blending manner.
+The emitted Gaussians are then rendered using a revised WSR paradigm.
+This pipeline conducts sorting at the cell level, establishing a coarse spatial order for patches of Gaussians.
+This strategy drastically reduces the computational burden of performing a radix sort on every individual Gaussian primitive (Sec.
+IV).
+Within each cell, Gaussians remain locally unsorted, making the order-independent WSR paradigm naturally suitable for their blending.
+This design eliminates popping artifacts, as the rendering result is locally invariant to the Gaussian order.
+Furthermore, by reintroducing physical constraints into the OIT-based renderer, a physically grounded transmittance decay is guaranteed, as shown in Fig.
+5, which removes the transparency artifacts inherent in the vanilla OIT-based Gaussian Splatting [16] (Sec.
+V) as well as enables early termination in cell-level to accelerate rendering.
+IV Layer I: Cell Proxy for 3DGS
+While neural Gaussian frameworks [20, 21] have achieved remarkable reconstruction accuracy, they still rely on conventional $\alpha$-blending operations as the introduced proxy anchors are points without concrete geometric interpretation.
+We introduce ellipsoidal cells with explicit visible regions as the proxies.
+Based on this representation, a cell rasterization strategy is introduced, enabling efficient and lightweight rasterization.
+Figure 4: Comparison between rasterization of GS and cell.
+Unlike the original $\alpha$-blending which processes each Gaussian individually, our method performs rasterization and early termination at the coarser cell level, where each cell encapsulates multiple Gaussians.
+The order confusion caused by overlap among cells is discussed in Sec.
+V-A.
 IV-A Cell Proxy
-Unlike anchors, which are points lacking geometric structure, each cell is represented as an ellipsoid that contains several neural 3D Gaussians whose physical properties are dynamically decoded from a trainable feature.
-Initialization begins with SfM points $\mathbf{P}\in\mathbb{R}^{N\times 3}$, from which we generate the initial cells, each centered at a point in accordance with the method described in [17].
-Every cell is parameterized by physical structure scales $\bm{S}_{n}$ and and quaternions $\bm{Q}_{n}$, defining its visible region.
+Unlike anchors, which are points lacking geometric structure, each cell is represented as an ellipsoid that covers several neural 3D Gaussians whose physical properties are dynamically decoded from a trainable feature.
+Initialization begins with SfM points $\mathbf{P}\in\mathbb{R}^{N\times 3}$.
+We generate the initial cells from these points, centering each one in accordance with the method described in [21].
+Every cell is parameterized by position $\bm{x}_{n}$, scales $\bm{S}_{n}$, quaternions $\bm{Q}_{n}$, defining its visible region, and a calibration scalar $v_{n}$ (discussed in Sec.
+V-A).
 In the dual-layered hierarchy, each cell is further associated with a latent feature vector $\bm{f}_{n}$ that encodes local scene information, and emits $K$ neural Gaussians as described in Eq.
-4.
-The attributes of these $K$ neural Gaussians—including color $\bm{c}_{n,k}$, scales $\bm{s}_{n,k}$ and quaternions $\bm{q}_{n,k}$—are decoded from the latent feature $\bm{f}_{n}$, along with the viewing distance $\delta_{n}$ and viewing direction $\bm{d}_{n}$, via shared scene-wise MLPs.
+1.
+The attributes of these $K$ neural Gaussians, including color $\bm{c}_{n,k}$, scales ratio $\mathcal{O}_{n,k}^{\text{scale}}$ and quaternions $\bm{q}_{n,k}$, are decoded from the latent feature $\bm{f}_{n}$, along with the viewing distance $\delta_{n}$ and viewing direction $\bm{d}_{n}$, via shared scene-wise MLPs.
+$\{\bm{c}^{n,1},...,\bm{c}^{n,K}\}=F_{c}(f_{n},\delta_{n},d_{n}),$ (8)
+$\{\mathcal{O}^{\text{scale}}_{n,1},...,\mathcal{O}^{\text{scale}}_{n,K}\}=F_{s}(f_{n},\delta_{n},d_{n}),$ (9)
+$\{\bm{q}_{n,1},...,\bm{q}_{n,K}\}=F_{q}(f_{n},\delta_{n},d_{n}).$ (10)
 To ensure that each Gaussian center lies within the visible region of its corresponding cell, we constrain the offset parameter $\mathcal{O}_{n,k}^{\text{pos}}$ to the range [1,1][-1,1].
-Additionally, to align the collective regions of the emitted neural Gaussians with the visible extent of the cell, we modify the encoding of the scale $\bm{s}_{n,k}$ and quaternions $\bm{q}_{n,k}$ attributes, thereby directly linking the geometric properties of the cell and its Gaussians
-$\{\mathcal{O}^{\text{scale}}_{n,1},...,\mathcal{O}^{\text{scale}}_{n,K}\}=F_{s}(f_{n},\delta_{n},d_{n}),$ (9)
-$\{\bm{q}_{n,1},...,\bm{q}_{n,K}\}=F_{q}(f_{n},\delta_{n},d_{n}),$ (10)
-where the output of $F_{s}$ and $F_{q}$ are activated by sigmoid and normalization respectively.
-Furthermore, the overall geometric configuration of each cell is explicitly associated with that of its constituent Gaussians through
+We propose that cells are likely to emit Gaussians with similar geometric properties.
+So scales of Gaussians are derived under the supervision of $\bm{S}_{n}$:
 $\bm{s}_{n,k}=\mathcal{O}_{n,k}^{\text{scale}}*\bm{S}_{n},$ (11)
-$\bm{Q}_{n}=\text{norm}(\sum_{k}^{K}\bm{q}_{n,k}*\sigma_{n,k}),$ (12)
-To restrict that the Gaussian regions fully locate inside the cell region, the maximum of $\mathcal{O}_{n,k}^{\text{scale}}$ is clamped by
+where the output of $F_{s}$ is activated by sigmoid to restrict the ratio $\mathcal{O}_{n,k}^{\text{scale}}$ to the range $(0,1)$.
+Based on Eq.11, the scales of cells can be directly optimized.
+While the quaternions unable to optimize in the similar way, as they denote the orientation of the clustered Gaussians.
+To align regions of cell and Gaussians, the cell quaternions are updated as the weighted average of the associated Gaussians’ orientation
+$\bm{Q}_{n}=\text{norm}(\sum_{k}^{K}\bm{q}_{n,k}*\sigma_{n,k}).$ (12)
+Additionally, to restrict that the Gaussian regions fully locate inside the cell region, the maximum of $\mathcal{O}_{n,k}^{\text{scale}}$ is further clamped by
 […]
@@
 Leveraging this property, the cell projections can effectively guide the Gaussian blending process.
-IV-B Cell Searching Rasterization
-In the original 3DGS framework [7], views are rendered by rasterizing 3D Gaussians onto the image plane through a Gaussian search rasterization strategy.
-In contrast, the anchors introduced in [16], which form the proxies of the dual-layered hierarchy, do not directly participate in the rendering process; instead, rendering remains confined strictly to the Gaussian level.
-This results in a distribution mismatch between the two hierarchy levels, i.e., between anchors and 3D Gaussians.
-Building on our proposed geometric cell, we introduce a cell search rasterization technique in which cells, rather than individual Gaussians, serve as the primary rasterization units.
+IV-B Cell Rasterization
+Conventional $\alpha$-blending renders views by rasterizing 3D Gaussians onto the image plane via splatting, a process that necessitates a global Gaussian traversal and sort.
+We fundamentally reformulate this pipeline by introducing a cell-level rasterization technique, where ellipsoidal cells, rather than individual Gaussians, serve as the primary rendering primitives.
 As illustrated in Fig.
-3, we perform rasterization at the cell level during the pre-processing stage, which reduces memory access overhead.
-Specifically, each cell defines a visible region, and only those whose regions intersect with the viewing frustum are recorded and subsequently decoded into neural 3D Gaussians for the subsequent $\alpha$-blending stage.
-This strategy ensures that the distribution of cells and 3D Gaussians are naturally aligned across the two hierarchy levels, and restricts rasterization exclusively to the cell level.
-Following the tile based paradigm, cells are assigned to individual tiles, and radix sort is applied to the cell sequence as necessary.
-Because each cell encapsulates a local cluster of Gaussians, the total number of cells is substantially less than the number of individual Gaussians.
-As a result, both memory consumption and computational overhead during rasterization are significantly reduced.
+4, this approach significantly reduces the number of primitives handled.
+Specifically, each cell defines a bounded visible region; only cells intersecting the viewing frustum are selected and decoded into neural 3D Gaussians.
+This proxy-guided design inherently ensures spatial alignment between the cell distribution and the underlying Gaussians, confining computationally intensive rasterization operations exclusively to the coarser cell level.
+Operating within a tile-based paradigm, visible cells are assigned to corresponding tiles, and a radix sort is applied to the cell sequence.
+Since the number of cells is orders of magnitude smaller than that of Gaussians, this strategy yields substantial reductions in both memory consumption and computational overhead for the rasterization process.
 Figure 5: Comparison of three different transmittance: LC-WSR (Eq.
-8), vanilla 3DGS (Eq.
-3) and ours (Eq.
+7), vanilla $\alpha$-blending (Eq.
+4) and ours (Eq.
 16).
-The proposed transmittance is not strictly monotonically decreasing with increasing depth, indicating that the Gaussians are locally unordered.
-Nevertheless, our method achieves blending weights similar to those of vanilla 3DGS, without requiring an explicit Gaussian order.
+By reintroducing physical constraints, our model produces a transmittance curve similar to that of $\alpha$-blending, establishing it as a reliable criterion for early termination.
+Furthermore, the proposed transmittance is not strictly monotonic with depth, confirming the local order-independent nature of Gaussian blending in our framework.
 V Layer II: Physical Weighted Sum Rendering
-Duplex-GS reconstructs 3D scenes using ellipsoidal cells, enabling rasterization and sorting to be performed at the cell level, thereby significantly reducing rasterization overhead.
-However, without explicitly sorting the depths of individual Gaussians—a requirement in the original $\alpha$-blending operations grounded in physical modeling—the exact blending order becomes ambiguous.
-Sort-free Gaussians [12], an OIT extension of vanilla 3DGS, demonstrated that optimized components and parameters for high-fidelity view synthesis are governed solely by the mathematical constraints of the scene model.
-This weighted sum approach is well-suited for integrating 3DGS into modern GPU graphics pipelines and for deployment on resource-constrained devices.
-Nevertheless, abandoning physical constraints introduces “transparency” artifacts that degrade visual quality, and the absence of early termination further results in reduced rendering efficiency.
+Duplex-GS reconstructs 3D scenes using ellipsoidal cells, which enable rasterization and sorting at a coarse cell level to significantly reduce rasterization overhead.
+However, this approach forgoes the explicit per-Gaussian depth sorting required by the original $\alpha$-blending pipeline, which is grounded in physical light transport models.
+This introduces an ambiguity in the blending order.
+The Sort-free Gaussian method [16], an OIT extension of 3DGS, demonstrates that high-fidelity view synthesis can be achieved through a weighted sum of Gaussians, governed by mathematical rather than physical constraints.
+This paradigm is highly suitable for graphics rendering pipelines in modern GPU and resource-constrained devices.
+Nevertheless, abandoning the physical blending model introduces transparency artifacts that degrade visual quality, and the lack of early termination further compromises rendering efficiency.
 In this section, we integrate the OIT based rendering paradigm with a cell based Gaussian model, retrieving physical constraints for WSR based 3DGS.
 V-A Physical OIT Blending for 3DGS
-WSR [12] is a mathematically motivated rendering model, in which $\alpha_{i}$ in Eq.
-6 does not directly represent the meaningful opacity of a Gaussian and may even exceed one due to its lack of a physical basis.
-While this formulation facilitates the weighted sum rendering process, it inherently precludes the early-termination mechanisms used in conventional $\alpha$-blending.
-Obtaining the sorted order at the cell level is both memory and computationally efficient, since the number of cells is substantially fewer than that of the 3D Gaussians.
-Leveraging this, we redefine $\alpha$ as the opacity weighted by the Gaussian kernel, as specified in Eq.
-2, constraining its value to the physically meaningful range of $[0,1]$.
-Thanks to the approximate ordering provided by cell-level sorting—which adequately reflects the relative relationships among the contained Gaussians—we are able to enforce physical constraints within the mathematical WSR framework, without the need for precise ordering of all Gaussians.
-As illustrated in Fig.
-3, the final rendered view $\bm{C}$ is computed according to the weighted sum rendering formula:
+Although the WSR paradigm [16] provides a mathematically-grounded rendering model, it lacks a physical foundation.
+Consequently, the blending weight $\alpha_{i}$ in Eq.
+5 does not correspond to a physically meaningful opacity and is unconstrained, often exceeding unity.
+This formulation not only introduces transparency artifacts but also precludes the use of early termination when blending opaque surfaces, limiting its efficiency.
+Within this coarse-ordered structure obtained from cell rasterization, we redefine $\alpha$ as an opacity weighted by the Gaussian kernel (Eq.
+3), constraining its value to the physically plausible range of $[0,1]$.
+The final rendered view $\bm{C}$ is computed according to the weighted sum rendering formula:
 $\bm{C}=\frac{\sum_{n=1}^{N}w_{n}\sum_{k=1}^{K}\alpha_{n,k}\bm{c}_{n,k}}{\sum_{n=1}^{N}w_{n}\sum_{k=1}^{K}\alpha_{n,k}},$ (14)
@@
 $w_{n}=v_{n}\cdot T_{n}^{\text{cell}},$ (15)
-where $T_{n}^{\text{cell}}$ denotes the cell-level transmittance, which is efficiently obtained by sorting the sparsely distributed cells and can be calculated as:
-$T_{n}^{\text{cell}}=\prod_{i=1}^{n-1}\prod_{k=1}^{K}(1-v_{i}\cdot\alpha_{i,k}).$ (16)
-Due to the non-uniform distribution and potential overlap of cells, $T_{i}^{\text{3DGS}}$ in Eq.
-3 may not precisely capture the transmittance of all emitted neural Gaussians.
-The introduction of the learnable parameter $v_{n}$ for each cell effectively compensates for this approximation.
-Three relevant paradigms of transmittance and the resulting Gaussians weights are visualized in Fig.
+where $T_{n}^{\text{cell}}$ denotes the cell-level transmittance, which is efficiently obtained by sorting the sparsely distributed cells, and it decays based on the accumulated opacity from preceding cells:
+$T_{n}^{\text{cell}}=\prod_{i=1}^{n-1}\prod_{k=1}^{K}(1-v_{i}\cdot\alpha_{i,k}),$ (16)
+where a learnable calibration scalar $v_{i}$ is introduced to address the ambiguity arising from non-uniform spatial distribution and potential inter-cell overlaps, as illustrated in Fig.
 4.
-LC-WSR simulates the decay trend of transmittance with a linear function w.r.t.
-Gaussian depth, introducing great number of Gaussians with tiny weights.
-Our method obtain similar weights compared with the vanilla 3DGS without the requirements of concrete order of individual Gaussians.
+Since cell occlusion relationships are view-dependent, we parameterize $v_{i}$ using spherical harmonic coefficients to capture its directional variation.
+The resulting Eq.
+14 shows that our framework produces model which is ordered outside cells and order-independent inside each cell.
+The opacity distribution in a 3D scene is typically sparse, implying that the accumulated transmittance is physically modeled by a step-function which suddenly changes encountering points with high opacity rather than a gradual decay.
+To validate this theory, we provide three transmittance paradigms and the resulting Gaussians weights in Fig.
+5.
+LC-WSR simulates transmittance with a linear function of Gaussian depth.
+However, this model is physically inconsistent, as it is derived from purely mathematical constraints.
+In contrast, our method produces a weight distribution similar to vanilla $\alpha$-blending without requiring a precise per-Gaussian depth order, which eliminates transparency artifacts for that the transmittance decays rapidly when encountered opaque object.
+Furthermore, the rendering of Gaussian primitives is locally order-independent, as the transmittance does not strictly decrease with depth.
+This property forms the cornerstone for the elimination of popping artifacts.
 V-B Early Termination for Cell Based 3DGS
-We propose to perform sorting based on cells, rather than individual Gaussians, thereby significantly reducing sorting overhead while retaining the hardware-friendly weighted sum rendering paradigm.
-As $T_{n}^{\text{cell}}$ decreases during the front-to-back rendering of cells, early termination can be efficiently incorporated to accelerate the rendering process and mitigate transparency artifacts.
-Specifically, $T_{n}^{\text{cell}}$, defined in Eq.
-16, approaches zero as cell-level transmittance accumulates and saturates, which provides a practical criterion for early termination: the blending process is halted when $T_{n}^{\text{cell}}<\epsilon$.
-Furthermore, in contrast to previous WSR methods for 3DGS, our approach restores explicit physical meaning to the opacity parameter $\sigma_{n,k}$, restricting it to the range $[0,1]$.
-This property facilitates pruning and regularization during training.
+As defined in Eq.
+16, the cell-level transmittance $T_{n}^{\text{cell}}$ decays towards zero during front-to-back rendering, eliminating transparency artifacts.
+This provides a natural mechanism for early termination: the blending process is halted in cell level once $T_{n}^{\text{cell}}$ falls below a predefined threshold $\epsilon$, as illustrated in Fig.
+4.
+This approach effectively accelerates rendering.
+Unlike prior WSR methods for 3DGS, our formulation restores explicit physical meaning to the opacity parameter $\sigma_{n,k}$ by constraining it to the physically plausible range of $[0,1]$, which means the empirical threshold can be directly transferred to our model.
+This enforcement of physical plausibility subsequently facilitates more effective pruning and regularization during the training process.
 VI Optimization Strategies for Cell Based Gaussian Splatting
-The proposed cell primitives are endowed with distinct physical attributes; however, these features are not directly involved in the differentiable blending process.
-Furthermore, advanced optimization strategies are introduced to systematically organize the evolving cell structures.
-VI-A Alignment of Cell and Gaussian Region
-The cell search rasterization inevitably leads to the introduction of redundant 3D Gaussians during the blending stage, due to the discrepancies that arise between the spatial intervals of cells and the emitted neural Gaussians.
-Furthermore, the efficiency of cell search rasterization is contingent upon the degree of coherence between the visible regions of a cell and its associated Gaussians.
-VI-A1 False positive skip
+The cell search rasterization inevitably introduces redundant 3D Gaussians during the blending stage, due to the discrepancies that arise between the spatial intervals of cells and the emitted neural Gaussians as shown in Fig.
+6a.
+As the efficiency of cell search rasterization is contingent upon the degree of coherence between the visible regions of a cell and its associated Gaussians, we provide several engineering-driven strategies to facilitate training of Duplex-GS.
+VI-A False Positive Skip
 Figure 6: Illustration of mismatch between cell and Gaussian.
@@
 (b) False positive regions are skipped to alleviate the speed degradation from the mismatch.
-TBR is widely adopted in modern GPUs and is also utilized in 3DGS rasterization.
-However, this approach often introduces a substantial number of redundant false-positive Gaussians, a problem that is further exacerbated in cell based Gaussian frameworks due to the spatial mismatch between cell regions and the actual coverage of Gaussians.
+Tile-based Rendering (TBR) is widely adopted in modern GPUs and is also utilized in 3DGS rasterization.
+However, this approach often introduces a substantial number of redundant false-positive Gaussians.
+This problem is further exacerbated in cell-based Gaussian frameworks due to the spatial mismatch between cell regions and the actual coverage of Gaussians.
 To mitigate the associated performance degradation, we introduce a pixel-level radius based culling strategy to eliminate unnecessary Gaussian kernel computations.
 Specifically, as illustrated in Fig.
-5, a lightweight axis-aligned bounding box (AABB) check is employed: Gaussians are excluded from the blending operation if the pixel lies outside their respective AABBs.
+6, a lightweight axis-aligned bounding box (AABB) check is employed: Gaussians are excluded from the blending operation if the pixel lies outside their respective AABBs.
 This efficient early-exit mechanism significantly reduces redundant computation and accelerates the overall rendering process.
-VI-A2 Geometric correction via neural Gaussians
+VI-B Geometric Correction via Neural Gaussians
 In addition to skipping false-positive Gaussians to improve efficiency, maximizing the overlap between each cell and its emitted neural Gaussians further reduces the number of redundant Gaussians.
@@
 12.
-Moreover, to ensure an even spatial distribution of Gaussians within each cell, the cell centers are periodically updated according to
+Moreover, to ensure an even spatial distribution of Gaussians within each cell, the cell centers are periodically updated using
 $\bm{x}_{n}=\frac{\sum_{k=1}^{K}\bm{\mu}_{n,k}}{K}.$ (17)
 As cell scales are differentiable via Eq. 11, the three-dimensional covariances and positions of first-layer cells are learned in accordance with the updates to the second-layer Gaussians.
-VI-A3 Cell position reset
-As discussed in Gaussian based models [7, 16, 17], the initial SfM point cloud often fails to comprehensively cover the entire scene, particularly in textureless or sparsely observed regions.
-Following the adaptive density control (ADC) policies employed in previous works [7], we utilize view-space positional gradients as a densification score to guide the growth of new cells.
-Specifically, additional cells are introduced at the centers of Gaussians that attain high densification scores.
-In contrast to the anchors proposed in [16, 17], which do not participate directly in the rendering process, our cells possess explicit physical attributes and independently encode local regions.
+VI-C Cell Position Reset
+Following adaptive density control [11], we grow new cells at Gaussians with high view-space positional gradients.
+Unlike abstract anchors [20, 21], our cells have explicit geometry.
 To decouple existing cells from newly introduced ones, the offsets of the significant Gaussians are reset to zero after cell growth, i.e., their positions are reinitialized to the centers of the corresponding cells.
-VI-B Perceptual Loss
-The loss function for optimizing 3D Gaussian Splatting (3DGS) is typically formulated as a combination of $\ell_{1}$ and SSIM [33] losses, which respectively capture pixel-level reconstruction error and structural similarity.
-Minimizing these terms promotes high signal-to-noise ratios but often results in images that lack high-frequency details and appear perceptually unsatisfying, particularly at higher resolutions.
-Given the effectiveness of perceptual loss functions in image generative models [34, 35], we additionally incorporate a perceptual loss term for training Gaussian based models.
+VI-D Perceptual Loss
+The loss function for optimizing 3DGS is typically formulated as a combination of $\ell_{1}$ and SSIM [56] losses, which respectively capture pixel-level reconstruction error and structural similarity.
+As proven in GAN-based models, these pixel-wise image differences results in solutions with high signal-to-noise ratios, which are perceptually rather smooth and less convincing [57, 58].
+We additionally incorporate a perceptual loss term for training Gaussian based models.
 The overall loss is defined as
@@
 LC-WSR2 [16] 27.21 0.800 0.219 1312.0 MB 23.22 0.832 0.186 672.7 MB 29.90 0.901 0.243 810.7 MB
+StocSplats3(16 SPP) [52] 26.25 0.714 0.351 - - - - - - - - -
 Ours $(K=10)$ 27.90 0.813 0.216 152.2 MB 24.26 0.867 0.150 124.8 MB 30.30 0.910 0.249 89.7 MB
@@
   • 2
-    The official implementation has not been released. Experiments are conducted using our own implementation, which is available at https://github.com/LiYukeee/sort-free-gs.
+    Experiments are conducted with our own implementation, which has been recognized by the author of [16]. Codes are available at https://github.com/LiYukeee/sort-free-gs.
+  • 3
+    Results from original paper (code unavailable). Missing data reflects source omissions.
 TABLE II: Quantitative comparisons on BungeeNeRF [63] and VR-NeRF [1] datasets.
@@
   • 2
-    The official implementation has not been released. Experiments are conducted using our own implementation, which is available at https://github.com/LiYukeee/sort-free-gs.
+    Experiments are conducted with our own implementation, which has been recognized by the author of [16]. Codes are available at https://github.com/LiYukeee/sort-free-gs.
+TABLE III: Quantitative comparisons on MatrixCity. Duplex-GS achieves highest accuracy across all metric in such large-scale aerial scenarios while maintaining competitive efficiency.
+Method MatrixCity
+$\text{PSNR}_{\uparrow}$ $\text{SSIM}_{\uparrow}$ $\text{LPIPS}_{\downarrow}$ $\text{\#GS/\#Proxy}_{\downarrow}$ $\text{FPS}_{\uparrow}$ $\text{Time~[h]}_{\downarrow}$
+3DGS1 [11] 27.03 0.808 0.310 10206 K 66 2.60
+Scaffold-GS [20] 26.82 0.811 0.288 4829 K 113 2.31
+Octree-GS [21] 25.18 0.738 0.397 4082 K 128 2.61
+Octree-GS* [21] 26.61 0.812 0.292 8010 K 79 -
+LC-WSR [16] 26.34 0.782 0.340 7841 K 34 3.86
+Ours $(K=5)$ 27.40 0.826 0.260 4165 K 119 2.21
+    Officially released pretrained model.
 VII-A Experimental Settings
 VII-A1 Datasets
-We conducted extensive experiments on 31 realistic scenes drawn from various datasets, following the configurations in 3DGS [7] and Scafolld-GS [16].
-Specifically, we evaluated 13 publicly released datasets, including nine from Mip-NeRF360 [36], two from Tanks and Temples [37] and two from DeepBlending [38].
-Additionally, we tested 10 further scenes—eight from BungeeNeRF [40] and two from VR-NeRF [1]—to assess the robustness of our approach in relatively large-scale environments.
-Scenes from BungeeNeRF feature multi-scale outdoor city scenes for validating view-adaptive rendering, while those from VR-NeRF provide intricate indoor scenes to evaluate detailed rendering capabilities.
+We conducted extensive experiments on scenes from multiple public datasets, adopting the standard configurations established by 3DGS [11].
+Our evaluation encompasses 13 standard benchmarks: nine scenes from Mip-NeRF360 [59], two from Tanks&Temples [60], and two from DeepBlending [61].
+To further assess robustness, we tested on 10 challenging scenes from BungeeNeRF [63], featuring multi-scale outdoor environments, and VR-NeRF [1], containing large-scale intricate interiors.
+Additionally, we included the large-scale aerial scene Block_All scene from MatrixCity [64], which comprises 5,621 training and 741 testing images covering an urban area of $2.7~\text{km}^{2}$.
+All experiments were performed at a resolution of 1K.
 VII-A2 Metrics
 We evaluate the reconstruction fidelity with PSNR, SSIM [56] and LPIPS [65], which are commonly used in image quality measurement.
-We also record the model size, frame per second (FPS) as indicators of storage overhead and rendering speed.
-Additionally, the consumptions of the radix sort are listed to address the effectiveness of our method in reducing the sorting overhead, which is a bottleneck for edge devices.
+We also record the model size, frame per second (FPS) and runtime breakdown as indicators of storage overhead and rendering speed tested on RTX 4090 GPU.
+Additionally, the consumptions of the radix sort are listed to address the effectiveness of our method in reducing the sorting overhead comparing with $\alpha$-blending, which is a bottleneck for edge devices.
+We also report FPS on a laptop with RTX 3060 GPU to verify the applicability of our method in mid-range devices.
 VII-A3 Baselines
-We compare our method with several SOTA approaches, including 3DGS, Mip-NeRF360, Scaffold-GS, Octree-GS, and LC-WSR.
-Specifically, LC-WSR serves as the baseline for OIT based Gaussian Splatting rendering, while the remaining methods employ the $\alpha$-blending paradigm used in the vanilla 3DGS.
+We compare our method with several SOTA approaches, including 3DGS, Mip-NeRF360, Scaffold-GS, Octree-GS, LC-WSR, and StocSplats.
+Among these, LC-WSR and StocSplats serve as representatives of the OIT based rendering paradigm for Gaussian splatting, while the remaining methods employ the conventional $\alpha$-blending pipeline used in vanilla 3DGS.
+The results for StocSplats are sourced directly from its original publication, which reports performance on the Mip-NeRF 360 dataset using an RTX 4090 GPU.
+As its training code and pre-trained models are not publicly available, we are unable to evaluate its performance on other benchmarks; the corresponding entries in our comparisons are therefore left blank.
 VII-A4 Implementation Details
-For the baseline models, we utilize the sum of the $\ell_{1}$ and SSIM losses as the loss function, with the weight $\lambda_{1}$ in Eq.
+All baseline models are optimized using a combined loss function comprising the $\ell_{1}$ and SSIM losses.
+The weighting factor $\lambda_{1}$ in Eq.
 18 is set to 0.2.
-For comparison, the perceptual loss is computed every 20 training iterations, with $\lambda_{2}$ set to $0.5$.
-For neural Gaussian based models (Scaffold-GS, Octree-GS, and the proposed Duplex-GS), experiments are conducted with encoding dimensions $K$ set to $5$ and $10$ respectively.
-For small-scale scenes containing hundreds of training images, training is performed for 40,000 iterations, with densification ending after 25,000 iterations.
-For large-scale scenes containing thousands of training images, training extends to 100,000 iterations, and densification concludes at 50,000 iterations.
-All experiments are conducted on a single NVIDIA RTX 4090 GPU.
+In contrast, a perceptual loss is computed every 20 training iterations with a weight $\lambda_{2}$ of 0.5.
+For neural Gaussian-based methods, including Scaffold-GS, Octree-GS, and our proposed Duplex-GS, experiments are conducted with encoding dimensions $K$ set to 5 and 10, respectively.
+Small-scale scenes are trained for 40,000 iterations, with densification concluding at iteration 25,000.
+Large-scale scenes are trained for 100,000 iterations, with densification ending at iteration 50,000.
 VII-B Results Analysis
-We first validate our methods by showing that it achieves the highest accuracy across several datasets. The results are shown in Tab. I and Tab. II.
+We first validate our method by showing that it achieves competitive accuracy across all datasets while maintaining high efficiency. The results are shown in Tab. I and Tab. II.
 Figure 7: Qualitative comparisons of Duplex-GS and competing baselines on Mip-NeRF360 [59] dataset.
@@
 I presents a comprehensive comparison of our method with several SOTA baselines on three real-world datasets.
-On the Mip-NeRF dataset, our approach achieves the highest PSNR when $K=10$, outperforming all baselines.
+On the Mip-NeRF360 dataset, our approach achieves the highest PSNR when $K=10$, outperforming all baselines.
 While Mip-Splatting and Octree-GS attain slightly higher SSIM values (0.816), our method yields competitive SSIM (0.813) and achieves improved perceptual quality, as evidenced by a lower LPIPS (0.216).
 Remarkably, our model also demonstrates strong storage efficiency, requiring significantly less memory than methods such as 3DGS and Mip-Splatting, both of which demand over 750 MB.
-On the Tanks&Temples and Deep Blending datasets, although the PSNR values of our approach are marginally lower, it delivers superior visual quality, attaining the highest SSIM and markedly lower LPIPS compared to other models.
-Notably, when compared to LC-WSR—the OIT variant of Gaussian Splatting—our method achieves better performance across all metrics and datasets, showing the effectiveness of the OIT paradigm in Gaussian Splatting on CUDA based high-end GPUs.
-To further assess our method under challenging scenarios involving multi-scale training views and large-scale datasets—common in real-world applications—we conduct experiments on the BungeeNeRF and VR-NeRF datasets (results shown in Table II).
+On the Tanks&Temples and DeepBlending datasets, although the PSNR values of our approach are marginally lower, it delivers superior visual quality, attaining the highest SSIM or lowest LPIPS compared to other models.
+Notably, when compared to LC-WSR, the OIT variant of Gaussian Splatting, our method achieves better performance across all metrics and datasets, showing the effectiveness of the OIT paradigm in Gaussian Splatting on CUDA based high-end GPUs.
+To further assess our method under challenging scenarios involving multi-scale training views and large-scale datasets, we conduct experiments on the BungeeNeRF, VR-NeRF and MatrixCity datasets (results shown in Tab.
+II and Tab.
+III).
 On BungeeNeRF, our approach achieves the highest PSNR and lowest LPIPS when $K=10$, and remains highly competitive with $K=5$, attaining the smallest model size among all baselines.
 The two scenes within VR-NeRF are large-scale, intricate indoor environments containing thousands of training images.
-Experimental results indicate that OIT based rendering consistently outperforms sort based $\alpha$-blending approaches for such cases, and our proposed method achieves the highest overall accuracy among all evaluated baselines.
+Experimental results indicate that OIT based rendering consistently outperforms sort-based $\alpha$-blending approaches for such cases.
+On the MatrixCity dataset, our method achieves the highest accuracy across all metrics (PSNR/SSIM/LPIPS) while maintaining competitive efficiency with the most efficient baseline (128 FPS vs.
+119 FPS) and the smallest training time.
 VII-B2 Qualitative Results
@@
 As illustrated in Fig.
-6 using examples from the Mip-NeRF360 dataset, our method eliminates the “transparency” artifacts observed in the previous OIT based Gaussian model, LC-WSR, while maintaining intricate scene details.
+7 and Fig.
+12 using examples from the Mip-NeRF360 dataset, our method eliminates the transparency artifacts observed in the previous OIT based LC-WSR, while maintaining intricate scene details.
 For multi-scale and large-scale scenarios, as shown in Fig.
-7, the proposed Duplex-GS successfully captures fine details, even in textureless and marginal regions.
-Additionally, our approach effectively removes “popping” artifacts, benefiting from the OIT blending kernel (see Fig.
-8).
+8, the proposed Duplex-GS successfully captures fine details, even in texture-less and marginal regions.
+Additionally, our approach effectively removes popping artifacts, benefiting from the OIT blending kernel (see Fig.
+11).
 Figure 8: Qualitative comparisons of Duplex-GS and competing baselines on BungeeNeRF [59] and VR-NeRF [1] datasets.
 the proposed Duplex-GS successfully captures fine details, even in textureless and marginal regions.
-TABLE III: Efficiency analysis of Duplex-GS and competing baselines.
+TABLE IV: Efficiency analysis of Duplex-GS and competing baselines, reporting the FPS at 1K resolution on RTX 4090 GPU.
 Method Mip-NeRF360 Tan.&Tem. DeepBlend. BungeeNeRF
@@
 LC-WSR [16] 77 89 114 31
-Ours 184 147 232 124
+StocSplats3 [52] 125 - - -
+Ours $(K=5)$ 184 147 232 124
+TABLE V: Runtime Breakdown (ms) on MipNeRF360.
+Method Stage bicycle bonsai counter flowers garden kitchen room stump treehill AVE
+LC-WSR rasterization 7.0177 2.7192 3.2817 3.8094 7.0134 4.5389 2.8045 4.1043 5.7666 4.5618
+blending 10.5629 4.2739 4.0905 4.9504 8.9972 6.0351 4.7305 5.2737 8.6236 6.5042
+sum 17.5806 6.9932 8.3722 8.7599 16.0106 10.5741 7.5350 9.3779 14.3903 11.0660
+Octree-GS decode 0.0027 0.0024 0.0030 0.0035 0.0038 0.0035 0.0018 0.0019 0.0021 0.0027
+rasterization 0.4232 0.3942 0.4723 0.3498 0.4280 0.4598 0.4497 0.3083 0.3823 0.4075
+sorting 0.6793 0.5985 0.8814 0.4757 0.6933 0.8888 0.7727 0.3011 0.4941 0.6427
+blending 1.3343 1.2240 1.4663 1.0070 1.4140 1.5561 1.5236 0.8118 1.0878 1.2695
+sum 2.4394 2.2191 2.8231 1.8367 2.5391 2.9082 2.7477 1.4230 1.9663 2.3225
+Ours decode 0.0028 0.0027 0.0028 0.0025 0.0032 0.0030 0.0023 0.0023 0.0023 0.0027
+rasterization 0.2562 0.2214 0.2387 0.2047 0.2200 0.2377 0.1944 0.2284 0.2580 0.2289
+sorting 0.3799 0.4169 0.6360 0.1576 0.2714 0.5857 0.3576 0.1465 0.1925 0.3493
+blending 1.7981 1.3834 1.8776 0.9951 1.6763 1.6698 2.1289 0.8722 1.3549 1.5285
+sum 2.4371 2.0243 2.7552 1.3600 2.1709 2.4963 2.6832 1.2493 1.8078 2.1093
 VII-B3 Efficiency Analysis
 Tab.
-III provides a comparison of rendering speeds, measured in FPS, between our method and several state-of-the-art baselines across four benchmark datasets.
+IV provides a comparison of rendering speeds, measured in FPS, between our method and several state-of-the-art baselines across four benchmark datasets.
 Our approach consistently achieves the highest or highly competitive FPS on all datasets.
@@
 These results highlight the efficiency and scalability of our approach, demonstrating real-time rendering capability and practical superiority in intricate and multi-scale scenarios.
+Tab.
+V presents a runtime breakdown to elucidate the mechanisms of our hybrid renderer.
+The analysis reveals that while LC-WSR benefits from hardware graphics pipeline compatibility, it performs suboptimally on TBR architectures due to redundant Gaussian primitives introduced in rendering (exemplified in Fig.
+5) and a lack of physical constraints.
+In contrast, our method achieves a significant acceleration in the rasterization and sorting stages.
+This gain is directly attributable to our coarse-grained processing of cell proxies, which drastically reduces the number of primitives handled compared to per-Gaussian methods.
+This efficiency introduces a trade-off: the blending phase exhibits increased latency.
+This is an expected consequence of our design, which retains sequential blending for cells and may process redundant Gaussians by operating on cell-based groups (of size $K$) rather than individual Gaussians.
+Despite this localized overhead, our approach achieves the highest overall rendering speed, demonstrating the effectiveness of the proposed hybrid paradigm.
 Radix sort is employed in 3DGS to determine the explicit ordering of Gaussian primitives, sorting $b$-digit numbers with $m$ possible values in $\mathcal{O}\left(b\cdot(n+m)\right)$ time and $\mathcal{O}(n+m)$ space.
@@
 To evaluate the effectiveness of our approach in reducing the computational and memory requirements of view-adaptive sorting, we record the sorting length and corresponding memory consumption during the radix sort procedure.
-The results, presented in Tab.
-IV, show that our method successfully mitigates the sorting overhead by introducing cell proxies, thereby demonstrating superior efficiency compared to existing approaches.
-TABLE IV: Comparisons about the overhead of radix sort stage. The proposed cell rasterization consistently reduce both the time and space consumptions compared with the vanilla 3DGS rasterization.
+Compared with methods adhering to conventional $\alpha$-blending paradigm, the results, presented in Tab.
+VI, show that our method successfully mitigates the sorting overhead by introducing cell proxies, thereby demonstrating superior efficiency.
+VII-B4 Hyperparameter Analysis
+We give analysis on the influence on the hyperparameter $K$, which controls the number of Gaussians decoded per cell.
+Since the opacity $\sigma$, decoded via $F_{\sigma}$ and activated by a Tanh function, must satisfy $\sigma>0$ for a Gaussian to be considered valid and included in rendering, it directly regulates the number of active primitives.
+The results show that while varying $K$ leads to only minor deviations in reconstruction accuracy, it substantially impacts computational efficiency.
+As illustrated in Fig.
+9, increasing $K$ results in more Gaussians being activated during rendering, which slows down the blending stage (Fig.
+9-a).
+Conversely, a higher $K$ reduces the number of proxies involved, thereby accelerating the rasterization process (Fig.
+9-b).
+This trade-off elucidates the performance characteristics of our hybrid renderer: the slower blending speed compared to $\alpha$-blending-based frameworks and significant acceleration over LC-WSR are attributable to the number of valid Gaussians (Fig.
+9-a), Furthermore, the results confirm that our cell-based framework exploits spatial sparsity more effectively than anchor-based approaches, as evidenced by the fewer proxies located within the same viewing frustum (Fig.
+9-b).
+Figure 9: Performance analysis with different hyperparameter $K$ on garden scenario from Mip-NeRF 360 dataset, which calculates metric (PSNR [dB]/FPS), the number of valid Gaussians and proxies inside a same viewing frustum during training.
+(a) The valid Gaussians is number of Gaussians with $\sigma>0$, indicating the efficiency of blending phase.
+(b) The involved proxies is the number of cells or anchors located inside the viewing frustum.
+The proposed cell is more sparse than anchors, resulting in acceleration in rasterization stage.
+TABLE VI: Comparisons about the overhead of radix sort stage. The proposed cell rasterization consistently reduce both the time and space consumptions compared with $\alpha$-blending based methods.
 Method Mip-NeRF360 BungeeNeRF
@@
 3DGS [11] 7,000,602 81.0 MB 21,798,876 255.0 MB
+Scaffold-GS [20] 4,588,785 53.69 MB 6,901,347 80.74 MB
+Octree-GS [21] 4,701,055 55.09 MB 5,480,410 64.14 MB
 Ours 3,217,845 37.7 MB 2,843,604 33.3 MB
 VII-C Ablation Studies
-VII-C1 Early Termination
-The introduction of the early termination mechanism is a key advantage of our method over previous approaches such as LC-WSR. To visualize the mechanism, we record the termination ratio with
+Figure 10: Visualization of ablation study. Quantitative results are shown in Tab. VII.
+VII-C1 Cell vs. Anchor
+The introduction of the ellipsoidal cell represents a key innovation, enabling efficient proxy rasterization and physical grounded WSR-based blending.
+In contrast, anchors adhere to the conventional $\alpha$-blending pipeline, which processes Gaussians individually.
+To validate the effectiveness of cells in proxy rasterization, we conduct an ablation study by replacing cells with anchor points while maintaining our hybrid rendering paradigm and retraining the models.
+As shown in Tab.
+VII, this anchor-based rasterization fails due to the lack of well-defined visible regions, confirming the necessity of our geometric cell structure.
+VII-C2 Cell-level Transmittance
+The physically grounded transmittance, shown in Eq.
+16, is the key innovation which eliminates transparency artifacts as well as enable early termination.
+We replace it with the conventional LC-WSR blending kernel in the proposed pipeline, which is marked as “w/o $T^{\text{cell}}$” in Tab.
+VII and Fig.
+10.
+The absence of physical constraints result in transparency artifacts and degraded accuracy.
+VII-C3 Early Termination (ET)
+The introduction of the early termination mechanism is a key advantage of our method over LC-WSR, which selectively omit computations that have negligible impact on the final rendered output with extremely small transmittance.
+To visualize the mechanism, we record the termination ratio with
 $r^{\text{ET}}=1-\frac{N^{\text{rendered}}}{N^{\text{val}}},$ (19)
 where $N^{\text{rendered}}$ denotes the number of Gaussians contributed to the final color, $N^{\text{val}}$ denotes that of the total interacted valid Gaussians along the ray.
+Since the proposed transmittance achieves a decay pattern similar to $\alpha$-blending (Fig.
+5), we maintain the same threshold value as used in vanilla 3DGS.
 The results are visualized in Fig.
-9.
-Our method demonstrates termination behavior which closely matches that of the vanilla 3DGS, whereas prior sort-free Gaussian Splatting methods lack this feature.
+13.
+Our method demonstrates termination behavior which closely matches that of $\alpha$-blending, whereas prior sort-free Gaussian Splatting methods lack this feature.
 To evaluate its effectiveness, we conduct an ablation study by disabling early termination during model training.
 As shown in Tab.
-V, enabling early termination leads to improvements in both accuracy and efficiency.
+VII, enabling early termination leads to improvements in both accuracy and efficiency.
 Figure 11: Illustration of “popping” artifacts removal by recording the differences between contiguous timestep.
+Figure 12: Illustration of “transparency” artifacts removal.
 Figure 13: Visualization of the early termination mechanism.
@@
 In contrast, Duplex-GS achieves termination results comparable to those of vanilla 3DGS, without requiring an explicit Gaussian order.
-TABLE V: Ablation experiments.
+TABLE VII: Ablation experiments. Visual examples are shown in Fig. 10.
 Method $\text{PSNR}_{\uparrow}$ $\text{SSIM}_{\uparrow}$ $\text{LPIPS}_{\downarrow}$ $\text{FPS}_{\uparrow}$
-w/o early termination 27.66 0.801 0.220 155
-w/o cell restriction 27.54 0.799 0.223 229
-w/o $v_{i}$ 27.59 0.801 0.218 174
+w/o cell 19.30 0.655 0.365 270
+w/o $T^{\text{cell}}$ 26.60 0.790 0.231 212
+w/o ET 27.66 0.801 0.220 155
+w/o alignment 27.54 0.799 0.223 229
+w/o $v_{n}$ 27.59 0.801 0.218 174
 Full method 27.74 0.802 0.218 184
-VII-C2 Cell Geometry Restriction
-Compared with anchor, cell has specific geometry, restricted by Eq.
+VII-C4 Alignment Restriction
+Unlike Gaussians whose geometry is directly optimized through blending operations, cells are updated guided by Gaussians and several constraints (Eq.
 12 and Eq.
-13, to align the regions between cell and the corresponding Gaussians.
-To show the effects of these conditions, we mute these restriction and the results are shown in Tab.
-V.
-The accuracy decreases as expected, however, the rendering speed is improved, which is regarded as the compromise between accuracy and efficiency.
-VII-C3 Transmittance Design
+13) are imposed to align coverage between cell and the corresponding Gaussians.
+To show the effectiveness of these conditions, we mute these restriction and the results are shown in Tab.
+VII.
+The accuracy decreases as expected, however, the rendering speed is improved, which offers a solution for resource-constrained platform for its significant acceleration while maintaining competitive visual quality.
+VII-C5 Order Confusion
 As the sorting results of cell level cannot reflect the exact order of Gaussians.
-we introduce learnable parameter $v_{i}$ to compensate the ambiguity.
-We train models without $v_{i}$ to show its effectiveness.
+we introduce learnable parameter $v_{n}$ to compensate the ambiguity.
+We train models without $v_{n}$ to show its effectiveness.
 The results are shown in Tab.
-V.
-By introducing $v_{i}$, the accuracy is improved and the efficiency is slightly improved due to the different amount of cell primitives.
+VII.
+By introducing $v_{n}$, the accuracy is improved as more degree-of-freedom is introduced to handle the overlap ambiguity among cells.
+TABLE VIII: Efficiency analysis of Duplex-GS and competing baselines on Mip-NeRF360 datasets, reporting FPS at 1K resolution on a laptop with RTX 3060 GPU .
+Method 3DGS LC-WSR Octree-GS Ours (KK=5) Ours (w/o alignment)
+FPS 42 21 47 49 72
+VII-D Performance on Mid-range GPUs
+We test rendering performance on a laptop with RTX 3060 GPU to validate the practicability of our method on mid-range devices.
+The results are shown in Tab.
+IV.
+Our method consistently achieves highest rendering speed while maintaining competitive visual quality.
 VIII Limitation and Conclusion
-In this work, we propose a dual-layered Gaussian model that effectively resolves the “popping” artifacts present in vanilla 3DGS and the “transparency” artifacts associated with OIT Gaussian Splatting.
-Leveraging cell search rasterization and an OIT weighted blending technique, our approach substantially reduces the overhead of global sorting.
-Implemented with CUDA, our method reconstructs intricate scene details while maintaining real-time rendering speed, thereby demonstrating the practicality and effectiveness of the OIT rendering paradigm for Gaussian Splatting models.
-Nevertheless, several limitations remain.
-Specifically, the geometric relationships between cells and Gaussians are not yet fully explored, leading to performance degradation due to intervals as illustrated in Fig.
-5.
-Additionally, while our current implementation is based on CUDA devices, it does not yet exploit intra-cell parallelism for Gaussian blending.
-Future work will aim to address these limitations, including efficient implementation on specialized hardware such as Field-Programmable Gate Array (FPGA) and Application-Specific Integrated Circuit (ASIC).
+This work presents a dual-hierarchy Gaussian model that effectively resolves popping artifacts inherent in vanilla $\alpha$-blending and the transparency artifacts of WSR.
+By introducing geometric proxy and a hybrid rendering paradigm, our approach substantially reduces the overhead associated with global sorting while achieving real-time performance as well as competitive visual quality.
+Several limitations suggest directions for future work.
+First, the geometric relationships between cells and their constituent Gaussians require further optimization, as imperfect spatial alignment can lead to performance degradation and redundant calculation, as illustrated in Fig.
+6.
+Second, the core WSR operation within each cell is inherently order-independent, presenting substantial, yet unexplored, parallel computing opportunities.
+Custom hardware accelerators based on FPGA or ASIC can capitalize on this parallelism to achieve higher performance and energy efficiency compared to general-purpose GPUs, thereby facilitating deployment on resource-constrained edge devices.